### v1

# Strong Normalization for the Safe Fragment of a Minimal Rewrite System: A Triple-Lexicographic Proof and a Conjecture on the Unprovability of Full Termination for Any Relational Operator-Only TRS

Abstract
We present a minimal operator-only term rewriting system with seven constructors and eight reduction rules. Our main contribution is a mechanically-verified proof of strong normalization for a guarded fragment using a novel triple-lexicographic measure combining a phase bit, multiset ordering (Dershowitz-Manna), and ordinal ranking. From strong normalization, we derive a certified normalizer with proven totality and soundness. Assuming local confluence (verified through critical pair analysis), Newman’s Lemma yields confluence and therefore unique normal forms for the safe fragment. We establish impossibility results showing that simpler measures, such as additive counters, polynomial interpretations, and single-bit flags, provably fail for rules with term duplication. The work demonstrates fundamental limitations in termination proving for self-referential systems. We state a conjecture: no relational operator-only TRS can have its full-system termination proved by internally definable methods. Here “relational” is equivalent to “capable of ordered computation”: systems with a recursor enabling iteration over successors, comparison, or sequential counting. Such recursors necessarily redistribute step arguments across recursive calls, defeating all additive termination measures. This structural limitation applies to any TRS expressive enough to encode ordered data. All theorems have been formally verified in a proof assistant. The formal development is available to program committee members and referees upon request for purposes of peer review.
1 Introduction
We develop a minimal operator-only rewrite calculus (KO7): the object language contains only constructors and operators with rewrite rules; there are no binders, types, external axioms, or semantic predicates. The rules are the semantics. Our goals are: (i) a clean, duplication-robust proof of strong normalization (SN); (ii) a certified normalizer that always returns a normal form; (iii) (optionally) unique normal forms via Newman’s Lemma under a local-confluence assumption; and (iv) a conjecture that no operator-only TRS capable of ordered computation can have its full-system termination proved by internally definable methods.
Scope. All formal results are established for a guarded safe subrelation. We do not claim a single global measure for the full unguarded relation. Moreover, we exhibit a precise negative: at the root peak $\mathrm{eqW}\,a\,a$ with $\kappa^{M}(a)=0$, local join fails in the full relation. Therefore the full relation is not locally confluent at that peak, hence not confluent. The designed remedy is to work in the safe fragment with guarded/context joiners (§6).
A second conceptual goal is to situate KO7 against results about fixed-target reachability in terminating TRSs. If we define an internal provability predicate by “$t$ reduces to $\top$”, then under SN the set $\{t\mid t\Rightarrow^{\ast}\top\}$ is decidable by normalization/backtracking. With confluence, decision reduces to normal-form equality. This explains why single-level Gödel encodings cannot coexist with globally terminating proof search; the right move is stratification.
Contributions.
(1) A duplication-robust SN proof for KO7 using a triple-lex measure with a multiset (DM) component and an MPO-style head precedence; (2) a total, proved-correct normalizer; (3) a guarded Newman module for the safe relation that yields confluence (hence unique normal forms) from SN + local confluence; (4) a decidability result for reachability under SN; (5) a catalog of impossibility results for additive and polynomial measures under duplication; (6) a formal verification of all results; (7) a conjecture on undecidable termination for operator-only systems within a fixed base theory.
Highlights (formalization summary).
  • SN (SafeStep) via triple lex: Formally proven using a lexicographic measure combining a $\delta$-phase bit, a Dershowitz–Manna multiset rank, and an ordinal.
  • Certified normalizer: A total and sound normalization function is defined by well-founded recursion.
  • Newman (SafeStep): Confluence is established via Newman’s Lemma using a verified local confluence property for the safe fragment.
  • Full Step caveat: We exhibit a specific peak ($\mathrm{eqW}\,a\,a$ with $\kappa^{M}(a)=0$) where local join fails, justifying the restriction to the safe subrelation.
  • Impossibility results: The failure of simpler additive and polynomial measures is formally witnessed by counterexamples.
2 Background: TRSs, SN, reachability, and Newman
We assume standard abstract reduction and term rewriting notions [BaaderNipkow98, Terese03]. A term is in normal form if no rule applies. A TRS is strongly normalizing (SN) if there are no infinite reductions. A relation is confluent if for any $t\Rightarrow^{\ast}u$ and $t\Rightarrow^{\ast}v$ there exists $w$ with $u\Rightarrow^{\ast}w$ and $v\Rightarrow^{\ast}w$. Local confluence requires this only for single-step forks. Newman’s Lemma asserts SN + local confluence $\Rightarrow$ confluence [Newman42], yielding unique normal forms.
Fixed-target (“small-term”) reachability in terminating TRSs has well-charted complexity: NP-complete for length-reducing systems (dropping to P under confluence), NExpTime/N2ExpTime for (linear) polynomial interpretations, and PSPACE for KBO-terminating TRSs [BaaderGiesl24]. Modularity holds in certain linear, non-collapsing combinations [CaronCoquide94], while even flat non-linear systems exhibit undecidable reachability and confluence [Mitsuhashi06]. Termination under duplication typically requires orders beyond plain sizes; a naive size can increase under duplicating rules. The cure is a multiset extension of a base order [DershowitzManna79] or (recursive) path orders [Dershowitz87].
3 The KO7 calculus
KO7 is a finite TRS over a small signature (7 constructors) with 8 rules (including an explicit case split on eqW). For concreteness we list the rule shapes below.
Rule Head Arity Shape Dup?
R1 merge 2 $\mathrm{merge}\;\mathrm{void}\;t\to t$ No
R2 merge 2 $\mathrm{merge}\;t\;\mathrm{void}\to t$ No
R3 merge 2 $\mathrm{merge}\;t\;t\to t$ No
[…]
R6 integrate 1 $\mathrm{integrate}\,(\mathrm{delta}\,t)\to\mathrm{void}$ No
R7 eqW 2 $\mathrm{eqW}\;a\;a\to\mathrm{void}$ No
R8 eqW 2 $\mathrm{eqW}\;a\;b\to\mathrm{integrate}\,(\mathrm{merge}\,a\,b)$ (if $a\neq b$) No
Table 1: KO7’s 8 rules. R7/R8 provide a complete case split on equality. R3 is collapsing. R5 redistributes $s$ to two RHS positions.
Tiny example (trace consequences).
Using the verified normalizer, we observe:
  • Integrate/delta: $\mathrm{integrate}(\delta\,t)\Rightarrow^{\ast}\mathrm{void}$ (verified).
  • Equality (meta-level consequence under confluence):
    • If $\mathrm{nf}(a)=\mathrm{nf}(b)$ then $\mathrm{eqW}\,a\,b\Rightarrow^{\ast}\mathrm{void}$.
    • Otherwise $\mathrm{eqW}\,a\,b\Rightarrow^{\ast}\mathrm{integrate}(\mathrm{merge}\,a\,b)$.
    Note: In the safe fragment, confluence ensures these outcomes are unique.
Note. We prove SN and Newman-based confluence for the safe fragment.
4 Strong normalization
We define a triple-lexicographic measure
$\mu^{3}(t):=(\delta\text{-flag}(t),\;\kappa^{M}(t),\;\mu_{\mathrm{ord}}(t))$
ordered by the lex product of: (i) a phase bit dropping on the successor recursion; (ii) a multiset of ranks $\kappa^{M}$ (Dershowitz–Manna) with an explicit precedence/status orienting redex >> pieces; and (iii) an ordinal payload $\mu_{\mathrm{ord}}$ for non-duplicating ties. Ordinal hazards are stated explicitly (right-addition is not strictly monotone; absorption $\alpha+\beta=\beta$ requires $\omega\leq\beta$). In duplicating branches (if extended to broader systems) we would use a compact MPO-style head precedence.
Step vs SafeStep (measure unification).
For the SafeStep relation we use a single unified triple lex order $(\delta,\kappa^{M},\mu_{\mathrm{ord}})$. For the full relation we only provide a disjunctive decrease certificate: each step is covered by either a KO7 lex drop or an MPO triple drop. We do not claim a single global well-founded measure for all full steps.
Theorem 1 (Per-step decrease (SafeStep)).
For every rule instance $t\Rightarrow t^{\prime}$ in the guarded SafeStep relation, we have $\mu^{3}(t^{\prime})<_{\mathrm{Lex}}\mu^{3}(t)$.
[…]
DM vs MPO on rules (explicit).
For merge-cancel, we use a DM multiset lift over a base order where $\mathrm{merge}\,t\,t$ strictly dominates $t$. For eqW, we use a compact MPO-leaning measure where the head precedence orients $\mathrm{eqW}\,a\,b$ strictly above $\mathrm{integrate}(\mathrm{merge}\,a\,b)$.
  • Base-order premise (merge-cancel): in $\mathrm{merge}\,t\,t\to t$, the RHS is strictly smaller than the LHS.
  • Base-order premise (eqW): in $\mathrm{eqW}\,a\,b\to\mathrm{integrate}(\mathrm{merge}\,a\,b)$, the RHS is strictly smaller under head precedence.
No-Go for constant bumps on $\kappa$ (generic duplicator).
Lemma 1 (No fixed +k+k or boolean flag orients a generic duplicator).
Let $\kappa$ be a max-depth-style counter and fix any $k\in\mathbb{N}$. For a duplicating rule of the shape $r(S)\to C[S,S]$ (one redex replaced by two occurrences of a subterm $S$), there exists an instance where $\kappa(\text{LHS}){+}k=\kappa(\text{RHS}){+}k$, so no strict lex drop occurs. The same holds for a boolean phase flag alone.
Proof sketch. Choose $S$ with $\kappa(S)\geq 1$ and let $\mathrm{base}$ bound the context. Then $\kappa(\text{LHS})=\mathrm{base}{+}1$ while $\kappa(\text{RHS})=\max(\kappa(C[S]),\kappa(C[S]))=\mathrm{base}{+}1$. Adding a fixed $k$ preserves equality; a single flag does not alter the tie. $\square$
Remark 1.
[…]
Duplication stress identity.
For any additive counter $\rho$ that counts a single removed redex and sums subpieces, generic duplicators satisfy
$\rho(\text{after})=\rho(\text{before})-1+\rho(S),$
so there is no strict drop when $\rho(S)\geq 1$. The robust fix uses DM/MPO: replace one element by a multiset of strictly smaller elements (DM), or use RPO/MPO with a precedence/status such that the LHS redex strictly dominates each RHS piece.
Corollary 1 (Strong normalization).
The guarded relation SafeStep for KO7 is strongly normalizing.
Genealogy of failures (why DM/MPO).
We record minimal counterpatterns that motivate the multiset/path components:
  • Pure ordinal ($\mu$) only: shape-blind bounds fail to separate nested $\delta$ from its context.
  • Additive bumps on $\kappa$: ties persist on duplicators (Lemma 1).
  • Additive counters $\rho$: by the identity $\rho(\text{after})=\rho(\text{before})-1+\rho(S)$ there is no strict drop when $\rho(S)\geq 1$.
  • Polynomial interpretations that rely on imported arithmetic or hard-coded constants: the duplicating rules force case splits that cannot be oriented without stepping outside the operator language, so these methods fail within the allowed internal toolbox.
The fix is DM/MPO: ensure each RHS piece is strictly smaller than the removed redex in a base order and lift via a multiset/path extension.
5 A certified normalizer
By well-founded recursion on $\mu^{3}$ we define a normalization function. We prove the following properties (separated to avoid overfull lines):
(Totality) $\forall t\,\exists n.~\mathrm{Normalize}(t)=n$
(Soundness) $\forall t.~\mathrm{normal}(\mathrm{Normalize}(t))\;\wedge\;t\Rightarrow^{\ast}\,\mathrm{Normalize}(t).$
These are exposed in the formalization. We make no efficiency claim: worst-case normalization cost follows the termination witness in play (cf. the small-term reachability bounds in [BaaderGiesl24]). In confluent systems, decision reduces to one normalization and an equality check.
6 Local confluence and Newman (guarded safe relation)
We discharge local confluence by joining the finite set of critical pairs (when present). Combining Cor. 1 with Newman’s Lemma [Newman42] yields:
Theorem 2 (Confluence and unique normal forms).
If a relation is strongly normalizing and locally confluent, then it is confluent. Hence every term reduces to a unique normal form.
Instantiation (SafeStep). Combining Cor. 1 (SN for SafeStep) with local-join lemmas yields confluence and unique normal forms for the safe fragment by Newman’s Lemma. In contrast, the full relation is not locally confluent at the root peak $\mathrm{eqW}\,a\,a$ under $\kappa^{M}(a)=0$; thus full confluence does not hold. The SafeStep relation restricts eqW to cases where arguments are guarded, preventing this divergence.
The star-star join proof follows the standard accessibility (Acc) recursion at the source, with a case split over star shapes (“head step + tail”) and composition via transitivity. The module also provides corollaries for uniqueness of normal forms and equality of normalizers under star.
Scope and Guarantees (formalization-accurate).
We work with a guarded safe subrelation SafeStep for which we prove:
  • SN (SafeStep). The KO7 triple measure $\mu^{3}$ strictly drops on every SafeStep. This yields a certified normalizer that is total and sound for the safe fragment.
  • Local Confluence (SafeStep). We provide local-join lemmas per root shape and context wrappers. Newman then yields confluence and unique NFs for the safe fragment.
  • Full Step per-rule decreases (Hybrid). For each kernel rule, there is a per-step decrease witnessed either by KO7’s $\delta/\kappa^{M}/\mu$ lex or by an MPO-leaning $\mu$-first triple. A uniform global aggregator for all of Step is left as future work.
Critical-pair coverage.
The following table covers the safe root configurations with explicit local-join lemmas. This is exhaustive for SafeStep except the reflexive $\mathrm{eqW}\,a\,a$ peak under $\kappa^{M}(a)=0$, which we show is not locally joinable at the root. We discharge many eqW cases via guarded/context wrappers.
Source Lemma
$\mathrm{integrate}(\delta\,t)$ localJoin_int_delta  (unique target void)
$\mathrm{merge}\;\mathrm{void}\;t$ localJoin_merge_void_left  (unique target $t$)
$\mathrm{merge}\;t\;\mathrm{void}$ localJoin_merge_void_right  (unique target $t$)
$\mathrm{merge}\;t\;t$ localJoin_merge_tt  (unique target $t$)
$\mathrm{rec}\,\Delta\,b\,s\,\mathrm{void}$ localJoin_rec_zero  (unique target $b$)
$\mathrm{rec}\,\Delta\,b\,s\,(\delta\,n)$ localJoin_rec_succ  (unique target $\mathrm{app}\,s\,(\mathrm{rec}\,\Delta\,b\,s\,n)$)
$\mathrm{eqW}\,a\,b,~a\neq b$ localJoin_eqW_ne  (unique target $\mathrm{integrate}(\mathrm{merge}\,a\,b)$)
$\mathrm{eqW}\,a\,a,~\kappa^{M}(a)=0$ not locally joinable at root
Guarded variants exclude spurious branches. Contextual wrappers lift root joins to context.
$\delta$-guard: definition and decidability.
Define the safe-phase predicate by $\delta\!\operatorname{-guard}(t)\iff\delta\!\operatorname{Flag}(t)=0$. Here $\delta\!\operatorname{Flag}:\mathrm{Term}\to\mathbb{N}$ is a structurally recursive function defined on terms in the artifact (tracking split parity), so the predicate is decidable. Facts: $\delta\!\operatorname{Flag}(\mathrm{eqW}\,a\,b)=0$; merge-void rules require $\delta\!\operatorname{Flag}(t)=0$.
[…]
$\mathrm{rec}\_\Delta\,b\,s\,(\delta\,n)$$\mathrm{app}\;s\;(\mathrm{rec}\_\Delta\,b\,s\,n)$$\delta$-flag drop
7 Impossibility results
We formally establish the failure of several simpler measure strategies, which necessitate the use of DM multiset orders or MPO. These negative results are verified in the formal development.
  • Additive bumps fail: No measure of the form $\kappa(t)+k$ strictly decreases on generic duplicating rules.
  • Bare flags fail: A single boolean flag is insufficient to orient rules that lift depth.
  • Polynomial interpretations fail: Any polynomial interpretation involving fixed constants requires external arithmetic axioms to “orient” the rule, violating the operator-only constraint.
Polynomial Impossibility.
Consider polynomial interpretations of the form $M(t)\in\mathbb{N}$ for a duplicating rule $f(x)\to g(x,x)$. A polynomial measure assigning $M(f(x))=M(x)+p$ requires $M(x)+p>2M(x)$ for strict decrease, implying $p>M(x)$. This fails for unbounded $M(x)$.
More generally: any polynomial proof for a system with duplicating rules must hardcode specific constants (e.g., $M(\mathrm{void})=2$) to satisfy inequalities like $2M(s)>M(s)+1$. Such proofs import external arithmetic properties and impose arbitrary values on operators, violating the principle that operators’ semantics should be defined solely by rewrite rules. Changing the hardcoded constant (e.g., $M(\mathrm{void})=1$) collapses the proof. This structural failure is why polynomial interpretations cannot orient generic duplicating rules without external axioms.
Internally definable measures.
We use a simple contract for internally defined termination measures to structure these negative results:
Definition 1 (Internally definable measure).
An internally definable measure for a type $\alpha$ consists of $(\beta,\;<\_{\beta},\;\mathsf{wf},\;m,\;\mathsf{ctxMono},\;\mathsf{piecesLt})$ where: $\beta$ is a base order carrier; $_{\beta}$ is a well-founded relation with witness $\mathsf{wf}$; $m:\alpha\to\beta$ is the measure; $\mathsf{ctxMono}$ expresses context compatibility; and $\mathsf{piecesLt}$ asserts that in each rule instance, every RHS piece is strictly smaller than the removed LHS redex w.r.t. $_{\beta}$.
8 Decidability of Reachability
Theorem 3 (Fixed-target reachability).
In a strongly normalizing TRS, the set $\{t\mid t\Rightarrow^{*}c\}$ for any constant $c$ is decidable via normalization.
This connects our work to classical results: if we could encode undecidable properties through reduction to constants, we would violate known theoretical limits. This motivates stratified approaches in proof assistants, where object-level and meta-level reasoning are carefully separated.
Assumptions (model).
We work with a finite first-order TRS over finite terms. The decision method is: compute a normal form (by SN) and check whether it is $\top$; with local confluence, this reduces to one normalization and an equality test.
Complexity context.
Small-term reachability in terminating TRSs ranges from NP (length-reducing) to NExpTime/N2ExpTime (polynomial interpretations) and PSPACE (KBO), with confluence lowering the length-reducing class to P [BaaderGiesl24]. This situates our “normalize and compare” decision procedure for KO7 within the established landscape.
Moreover, decidability can be modular for disjoint unions under left-linearity/non-collapsing assumptions [CaronCoquide94], while termination alone does not guarantee decidability: even flat non-linear TRSs have undecidable reachability/joinability/confluence [Mitsuhashi06].
9 Conjecture: Full Termination of Any Relational Operator-Only TRS
What “relational” means.
A relational operator-only TRS is, by definition, one capable of ordered computation. The terms are equivalent: “relational” = “capable of ordered computation.” Such systems can express iteration over successors, comparison of structure depths, or any form of sequential counting. The minimal signature for such capability includes a recursor, an operator that applies a step function iteratively across a successor-structured argument. Examples include primitive recursion over natural numbers, fold over lists, and any construct that steps through an ordered index. A TRS that cannot express such iteration (e.g., a flat pattern-matching system with no recursive structure) is not relational in this sense.
The critical property of relational systems is that their recursor redistributes its step argument across recursive calls. This redistribution is not optional: it is what makes ordered computation possible. But it also defeats all additive termination measures.
The structural barrier.
Any operator system capable of ordered computation requires a recursor. A recursor redistributes its step argument across recursive calls. This redistribution defeats additive termination measures: $M(\text{after})=M(\text{before})-1+M(s)$, yielding no strict drop when $M(s)\geq 1$.
Definition 2 (Internally definable measure).
An internally definable termination method for signature $\Sigma$ uses only: simplification orders with fixed precedence/status on $\Sigma$ (LPO/RPO/MPO), DM-multiset lifts of $\mathbb{N}$-valued ranks, and algebraic interpretations definable without importing external axioms. No encodings into external arithmetic, no borrowed logic, and no hard-coded constants outside the operator language.
Conjecture 1 (Full Termination Conjecture for Relational TRSs).
No relational operator-only TRS can have its full-system termination proved by internally definable methods. Specifically: let $R$ be an operator-only TRS with a recursor rule of the form
$\mathrm{rec}(b,s,\sigma(n))\to f(s,\mathrm{rec}(b,s,n)).$
No internally definable measure (Definition 2) proves termination of $R$ when the step argument $s$ is unrestricted. The structural redistribution of $s$ into both the function application and the recursive call defeats every additive, polynomial, or path-order measure that does not import external arithmetic or ordinal axioms. Any proof that relies on external axioms, imported arithmetic, or meta-level encodings is outside scope and does not count as an internal termination argument.
Scope of the conjecture.
The conjecture applies to any TRS expressive enough to encode ordered data (natural numbers, lists, trees with successor structure, or any type supporting iteration). Systems without such structure (purely ground, non-recursive, or lacking a step-iteration pattern) fall outside the conjecture’s scope. The claim is that once a system has enough structure for ordered computation, its full termination escapes internal proof methods.
Evidence.
In KO7, the rec-succ rule has the critical shape: the step argument $s$ appears twice on the RHS, and when $s$ contains nested redexes, every additive measure fails to decrease strictly. The redistribution identity $M(\text{after})=M(\text{before})-1+M(s)$ yields no strict drop when $M(s)\geq 1$. The guarded SafeStep fragment terminates because the $\delta$-phase bit restricts allowable step arguments; without such restriction, the full system admits unbounded recursive depth. The gap between internally definable ranking functions and what would be required cannot be closed without external methods (axioms, encodings, or borrowed arithmetic).
10 Formalization structure
The formal verification is implemented in a proof assistant. The project structure separates the kernel definitions, meta-theory proofs, and impossibility results:
  • Termination proofs: Establishes the per-rule decreases and strong normalization for the safe fragment using the triple-lexicographic measure.
  • Normalization: Defines the normalization function and proves its totality and soundness.
  • Confluence: Implements the Newman engine and the star-star join, relying on local join lemmas.
  • Impossibility results: Contains the verified counterexamples for additive measures and the proofs that simple measures fail under duplication.
All claimed results are formally proven without reliance on unproven postulates in the current build. The formal development is available to reviewers under appropriate confidentiality agreements.
Appendix: Addendum
KO7 Rules Table (7 constructors, 8 rules)
Rule Head Arity Shape Dup?
R1 merge 2 $\mathrm{merge}\;\mathrm{void}\;t\to t$ No
R2 merge 2 $\mathrm{merge}\;t\;\mathrm{void}\to t$ No
R3 merge 2 $\mathrm{merge}\;t\;t\to t$ No
[…]
R6 integrate 1 $\mathrm{integrate}\,(\mathrm{delta}\,t)\to\mathrm{void}$ No
R7 eqW 2 $\mathrm{eqW}\;a\;a\to\mathrm{void}$ No
R8 eqW 2 $\mathrm{eqW}\;a\;b\to\mathrm{integrate}\,(\mathrm{merge}\,a\,b)$ (if $a\neq b$) No
Table 2: KO7’s 8 rules. R7/R8 provide a complete case split on equality. R3 is collapsing (erases one copy). R5 redistributes $s$ to two positions on the RHS; see Remark below.
Triple-lexicographic measure and duplication handling (DM/MPO)
We order $\mu^{3}(t)=(\delta\text{-flag}(t),\,\kappa^{M}(t),\,\mu_{\mathrm{ord}}(t))$ by lex:
  • $\delta$-flag: phase bit dropping on the successor recursion branch.
  • Multiset $\kappa^{M}$: a Dershowitz–Manna multiset extension over a base precedence/status (or MPO head precedence) orienting redexes strictly above pieces; covers duplicators.
  • Ordinal $\mu_{\mathrm{ord}}$: resolves non-duplicating ties; right-addition is not strictly monotone; absorption $\alpha+\beta=\beta$ needs $\omega\leq\beta$.
Per-rule lemmas show every RHS component is strictly smaller than the removed redex in the base order.
Aggregation for $\kappa^{M}$ (union, not sum).
We emphasize that $\kappa^{M}$ aggregates via multiset union ($\cup$), not numeric addition. For duplicating rules, the multiset of piece-weights on the RHS is DM-smaller than the singleton multiset containing the LHS redex weight, yielding a strict drop in the $\kappa^{M}$ component. For non-duplicating rules, $\kappa^{M}$ ties by definitional equality and the ordinal $\mu_{\mathrm{ord}}$ resolves the branch. In particular, for unguarded instances of eqW-refl and merge-cancel we use a $\kappa$-branch: when $\kappa^{M}(a)\neq 0$ we obtain a left-lex drop via DM; when $\kappa^{M}(a)=0$ the $\kappa^{M}$ component ties by rfl and the strict decrease is witnessed in the $\mu_{\mathrm{ord}}$ coordinate (right branch of Prod.Lex).
Short witness snippets (toy duplication).
Toy rule: $\mathrm{pair}(\mathrm{s}\,x,\,y)\to\mathrm{pair}(x,\,\mathrm{pair}(y,\,y))$.
DM multiset on sizes. Let $S(x,y)=\mathrm{size}(\mathrm{pair}(\mathrm{s}\,x,\,y))=\mathrm{size}(x)+\mathrm{size}(y)+2$. Then $\mathrm{size}(x)<S(x,y)$ and $\mathrm{size}(y)<S(x,y)$. Use $X=\varnothing$, $Y=\{\mathrm{size}(x),\mathrm{size}(y)\}$, $Z=\{S(x,y)\}$ to conclude $Y<_{\mathrm{DM}}Z$.
MPO triple weight. $\mathrm{weight}(\mathrm{pair}\,a\,b)=(\mathrm{headRank}(a),\,\mathrm{size}(a),\,\mathrm{size}(b))$ with $\mathrm{headRank}(\mathrm{s}\,\_)=2$, else $1$. If $x$ is unit/pair then first components decrease (1¡2). If $x=\mathrm{s}\,t$ then tie on 2; the second components satisfy $\mathrm{size}(t)+1<\mathrm{size}(t)+2$.
[…]
Figure 1: Triple-lex measure components. Duplicators decrease via $\kappa^{M}$; non-duplicating ties via $\mu_{\mathrm{ord}}$.
Newman scope (guarded safe relation)
SN + local confluence implies confluence; in the artifact this is instantiated for the safe relation via an Acc-based star–star join. Scope note: all confluence statements and their Newman instantiations are for SafeStep only. Full Step is not locally joinable at root for eqW a a with $\kappa^{M}(a)=0$; accordingly we do not claim full-step confluence.
Module map
The verification logic is distributed across key modules for SN, normalizer, and confluence engine. Local confluence lemmas are proved separately. Note: context closure and local-join lemmas are explicitly instantiated for the SafeStep fragment (SafeStepCtx) rather than the full relation.
[…]
Rule Base component Precedence/Status Witness Source
merge void left $\mu$ (ordinal) Theorem 4.1 Termination
merge void right $\mu$ (ordinal) Theorem 4.2 Termination
merge cancel DM on $\kappa^{M}$ (redex >> pieces) head precedence Theorem 4.3 Termination
rec zero DM on $\kappa^{M}$ rec >> pieces Theorem 4.4 Termination
rec succ $\delta$ phase bit (1 $\to$ 0) Theorem 4.5 Termination
eqW refl MPO ($\mu$-first triple) head precedence Theorem 4.6 Termination
eqW diff MPO ($\mu$-first triple) head precedence Theorem 4.7 Termination
integrate (delta t) $\mu$ (unique target) Theorem 4.8 Termination
toy duplication DM on size s >> pair Lemma 7.1 Impossibility
toy duplication MPO triple headRank(s) >> headRank(pair) Lemma 7.2 Impossibility
Table 3: Per-rule orientation summary.
Parenthetical aliases: for substitution convenience we expose simp forms alongside the main rec lemmas.
Local-join report (critical peaks and unique-target cases)
Shape Join lemma name Notes
merge void void localJoin_merge_void_void both branches to void
merge t t localJoin_merge_tt joins at $t$ (context-lifted)
integrate (delta t) localJoin_int_delta unique target void
[…]
eqW a b, a $\neq$ b localJoin_eqW_ne only diff applies (unique target)
eqW a a, guard kappaM a $\neq$ 0 localJoin_eqW_refl_guard_ne refl blocked; only diff
merge a b (no void, a $\neq$ b) localJoin_merge_no_void_neq no root step; vacuous
integrate non-delta localJoin_integrate_non_delta no root step; vacuous
[…]
Table 4: Local-join witnesses for the safe relation (root).
Ctx-local join wrappers
When needed in context, the root local-join lemmas lift to context via wrappers:
Shape Ctx wrapper lemma Notes
merge a b (no void, a $\neq$ b) localJoin_ctx_merge_no_void_neq from root ‘merge’ lemma
eqW a b, a $\neq$ b localJoin_ctx_eqW_ne only diff applies
eqW a a, kappaM guard localJoin_ctx_eqW_refl_guard_ne refl blocked
[…]
eqW (delta n) (delta n) localJoin_ctx_eqW_refl_when_a_is_delta literal delta case
Table 5: Ctx-local join wrappers.
Appendix: Orientation Strategy
This summarizes the exact orientation strategy realized in the proofs (no global RPO required):
  • Non-duplicating rules use a $\mu$-first lex measure: triples $(\mu,\mathrm{sizeMPO},\delta)$ decrease directly via lemmas.
  • The rule rec_zero uses a DM multiset drop on $\kappa^{M}$, wrapped by drop_R_rec_zero.
  • The rule rec_succ drops the $\delta$ phase bit (1 $\to$ 0) via drop_R_rec_succ.
  • Merge-cancel decreases via DM on $\kappa^{M}$ under a base precedence where the redex strictly dominates its pieces.
  • Conceptual precedence (if one prefers an MPO reading) aligns with rec >> merge >> app >> integrate >> eqW >> delta >> void, but the formal proofs do not require a global precedence declaration.
Formal anchors:
  • Merge void left/right: $\mu$-first or DM tie plus $\mu$.
  • Merge cancel: DM multiset drop on $\kappa^{M}$.
  • Rec zero: drop_R_rec_zero with dm_drop_R_rec_zero for the inner DM piece.
  • Rec succ: drop_R_rec_succ via the $\delta$ phase bit (1 $\to$ 0).
  • EqW refl/diff: $\mu$-first MPO triple.
  • Integrate delta: unique target void.
Appendix A Related work
We follow standard expositions in Baader-Nipkow and Terese [BaaderNipkow98, Terese03]. Termination under duplication relies on DM multiset orders [DershowitzManna79] and (recursive) path orders [Dershowitz87]. Newman’s Lemma is used in its classic ARS form [Newman42]. For independence via termination, see Goodstein and Kirby-Paris and ordinal tools due to Buchholz [Goodstein44, KirbyParis82, Buchholz86]. For fixed-target reachability complexity and modularity results, see [BaaderGiesl24, CaronCoquide94]; for undecidability at shallow depth in non-linear systems, see [Mitsuhashi06].
Appendix B Conclusion
We gave an operator-only kernel and proved strong normalization and confluence for the guarded safe subrelation, yielding a certified normalizer and unique normal forms via Newman’s Lemma. We also established an explicit non-local-join witness at the eqW reflexive peak (under $\kappa^{M}=0$), blocking confluence for the full relation; this motivates the designed SafeStep fragment. A decidability result explains why single-level Gödel encodings clash with global termination. We proved a simple impossibility for additive bumps under duplication and certified DM/MPO orientations. Future work: complete a stratified meta-layer; develop arithmetic encodings within KO7; and map the exact boundary between internally definable measures and external ordinal strength.

### v2

@@
 Abstract
-We present a minimal operator-only term rewriting system with seven constructors and eight reduction rules.
-Our main contribution is a mechanically-verified proof of strong normalization for a guarded fragment using a novel triple-lexicographic measure combining a phase bit, multiset ordering (Dershowitz-Manna), and ordinal ranking.
-From strong normalization, we derive a certified normalizer with proven totality and soundness.
+This paper presents a minimal operator-only term rewriting system with seven constructors and eight reduction rules.
+The main contribution is a mechanically-verified proof of strong normalization for a guarded fragment using a novel triple-lexicographic measure combining a phase bit, multiset ordering (Dershowitz-Manna), and ordinal ranking.
+From strong normalization, a certified normalizer with proven totality and soundness is derived.
 Assuming local confluence (verified through critical pair analysis), Newman’s Lemma yields confluence and therefore unique normal forms for the safe fragment.
-We establish impossibility results showing that simpler measures, such as additive counters, polynomial interpretations, and single-bit flags, provably fail for rules with term duplication.
+Impossibility results showing that simpler measures, such as additive counters, polynomial interpretations, and single-bit flags, provably fail for rules with term duplication, are established.
 The work demonstrates fundamental limitations in termination proving for self-referential systems.
-We state a conjecture: no relational operator-only TRS can have its full-system termination proved by internally definable methods.
+A conjecture is stated: no relational operator-only TRS can have its full-system termination proved by internally definable methods.
 Here “relational” is equivalent to “capable of ordered computation”: systems with a recursor enabling iteration over successors, comparison, or sequential counting.
@@
 All theorems have been formally verified in a proof assistant.
-The formal development is available to program committee members and referees upon request for purposes of peer review.
+The Lean formalization is available at https://github.com/MosesRahnama/OperatorKO7.
 1 Introduction
-We develop a minimal operator-only rewrite calculus (KO7): the object language contains only constructors and operators with rewrite rules; there are no binders, types, external axioms, or semantic predicates.
-The rules are the semantics.
-Our goals are: (i) a clean, duplication-robust proof of strong normalization (SN); (ii) a certified normalizer that always returns a normal form; (iii) (optionally) unique normal forms via Newman’s Lemma under a local-confluence assumption; and (iv) a conjecture that no operator-only TRS capable of ordered computation can have its full-system termination proved by internally definable methods.
+This paper presents a minimal operator-only rewrite calculus (KO7) where the object language contains only constructors and operators with rewrite rules.
+There are no binders, types, external axioms, or semantic predicates; the rules are the semantics.
+The primary goals are: (i) a clean, duplication-robust proof of strong normalization (SN); (ii) a certified normalizer that always returns a normal form; (iii) unique normal forms via Newman’s Lemma under a local-confluence assumption; and (iv) a conjecture regarding the unprovability of termination for relational operator-only systems using internal methods.
 Scope.
-All formal results are established for a guarded safe subrelation.
-We do not claim a single global measure for the full unguarded relation.
-Moreover, we exhibit a precise negative: at the root peak $\mathrm{eqW}\,a\,a$ with $\kappa^{M}(a)=0$, local join fails in the full relation.
-Therefore the full relation is not locally confluent at that peak, hence not confluent.
-The designed remedy is to work in the safe fragment with guarded/context joiners (§6).
+The formal results apply to a guarded safe subrelation.
+This restriction is necessary because the full relation is not locally confluent at the root peak $\mathrm{eqW}\,a\,a$ when $\kappa^{M}(a)=0$.
+The safe fragment employs guarded/context joiners to ensure unique normal forms (§6).
 A second conceptual goal is to situate KO7 against results about fixed-target reachability in terminating TRSs.
-If we define an internal provability predicate by “$t$ reduces to $\top$”, then under SN the set $\{t\mid t\Rightarrow^{\ast}\top\}$ is decidable by normalization/backtracking.
+If an internal provability predicate is defined by “$t$ reduces to $\top$”, then under SN the set decidability of $\{t\mid t\to^{*}\top\}$ is reducible to normalization.
 With confluence, decision reduces to normal-form equality.
-This explains why single-level Gödel encodings cannot coexist with globally terminating proof search; the right move is stratification.
+This explains why single-level Gödel encodings cannot coexist with globally terminating proof search; stratification is required.
 Contributions.
-(1) A duplication-robust SN proof for KO7 using a triple-lex measure with a multiset (DM) component and an MPO-style head precedence; (2) a total, proved-correct normalizer; (3) a guarded Newman module for the safe relation that yields confluence (hence unique normal forms) from SN + local confluence; (4) a decidability result for reachability under SN; (5) a catalog of impossibility results for additive and polynomial measures under duplication; (6) a formal verification of all results; (7) a conjecture on undecidable termination for operator-only systems within a fixed base theory.
+Strong normalization is proven for a guarded fragment of KO7 using a novel triple-lexicographic measure that combines a phase bit, multiset ordering (Dershowitz-Manna), and ordinal ranking.
+This result supports a certified normalizer with proven totality and soundness.
+Furthermore, confluence for the safe fragment is established via Newman’s Lemma given verified local confluence.
+The work also includes formally verified impossibility results for additive and polynomial measures under duplication, demonstrating fundamental limitations in termination proving for self-referential systems.
 Highlights (formalization summary).
-  • SN (SafeStep) via triple lex: Formally proven using a lexicographic measure combining a $\delta$-phase bit, a Dershowitz–Manna multiset rank, and an ordinal.
+  • SN (SafeStep) via triple lex: Formally proven using a lexicographic measure combining a $\delta$-phase bit, a Dershowitz-Manna multiset rank, and an ordinal.
   • Certified normalizer: A total and sound normalization function is defined by well-founded recursion.
   • Newman (SafeStep): Confluence is established via Newman’s Lemma using a verified local confluence property for the safe fragment.
-  • Full Step caveat: We exhibit a specific peak ($\mathrm{eqW}\,a\,a$ with $\kappa^{M}(a)=0$) where local join fails, justifying the restriction to the safe subrelation.
+  • Full Step caveat: A specific peak ($\mathrm{eqW}\,a\,a$ with $\kappa^{M}(a)=0$) where local join fails is exhibited, justifying the restriction to the safe subrelation.
   • Impossibility results: The failure of simpler additive and polynomial measures is formally witnessed by counterexamples.
 2 Background: TRSs, SN, reachability, and Newman
-We assume standard abstract reduction and term rewriting notions [BaaderNipkow98, Terese03].
+Standard abstract reduction and term rewriting notions apply [2, 11].
 A term is in normal form if no rule applies.
@@
 Local confluence requires this only for single-step forks.
-Newman’s Lemma asserts SN + local confluence $\Rightarrow$ confluence [Newman42], yielding unique normal forms.
-Fixed-target (“small-term”) reachability in terminating TRSs has well-charted complexity: NP-complete for length-reducing systems (dropping to P under confluence), NExpTime/N2ExpTime for (linear) polynomial interpretations, and PSPACE for KBO-terminating TRSs [BaaderGiesl24].
-Modularity holds in certain linear, non-collapsing combinations [CaronCoquide94], while even flat non-linear systems exhibit undecidable reachability and confluence [Mitsuhashi06].
+Newman’s Lemma asserts SN + local confluence $\Rightarrow$ confluence [10], yielding unique normal forms.
+Fixed-target (“small-term”) reachability in terminating TRSs has well-charted complexity: NP-complete for length-reducing systems (dropping to P under confluence), NExpTime/N2ExpTime for (linear) polynomial interpretations, and PSPACE for KBO-terminating TRSs [1].
+Modularity holds in certain linear, non-collapsing combinations [4], while even flat non-linear systems exhibit undecidable reachability and confluence [9].
 Termination under duplication typically requires orders beyond plain sizes; a naive size can increase under duplicating rules.
-The cure is a multiset extension of a base order [DershowitzManna79] or (recursive) path orders [Dershowitz87].
+The cure is a multiset extension of a base order [5] or (recursive) path orders [6].
 3 The KO7 calculus
-KO7 is a finite TRS over a small signature (7 constructors) with 8 rules (including an explicit case split on eqW). For concreteness we list the rule shapes below.
+KO7 is a finite TRS over a small signature (7 constructors) with 8 rules (including the equality-witness operator eqW). For concreteness we list the kernel rule shapes below.
 Rule Head Arity Shape Dup?
-R1 merge 2 $\mathrm{merge}\;\mathrm{void}\;t\to t$ No
-R2 merge 2 $\mathrm{merge}\;t\;\mathrm{void}\to t$ No
-R3 merge 2 $\mathrm{merge}\;t\;t\to t$ No
+R_merge_void_left merge 2 $\mathrm{merge}\;\mathrm{void}\;t\to t$ No
+R_merge_void_right merge 2 $\mathrm{merge}\;t\;\mathrm{void}\to t$ No
+R_merge_cancel merge 2 $\mathrm{merge}\;t\;t\to t$ No
 […]
 […]
-R6 integrate 1 $\mathrm{integrate}\,(\mathrm{delta}\,t)\to\mathrm{void}$ No
-R7 eqW 2 $\mathrm{eqW}\;a\;a\to\mathrm{void}$ No
-R8 eqW 2 $\mathrm{eqW}\;a\;b\to\mathrm{integrate}\,(\mathrm{merge}\,a\,b)$ (if $a\neq b$) No
-Table 1: KO7’s 8 rules. R7/R8 provide a complete case split on equality. R3 is collapsing. R5 redistributes $s$ to two RHS positions.
+R_int_delta integrate 1 $\mathrm{integrate}\,(\mathrm{delta}\,t)\to\mathrm{void}$ No
+R_eq_refl eqW 2 $\mathrm{eqW}\;a\;a\to\mathrm{void}$ No
+R_eq_diff eqW 2 $\mathrm{eqW}\;a\;b\to\mathrm{integrate}\,(\mathrm{merge}\,a\,b)$ No
+Table 1: KO7’s 8 kernel rules.
+In the full kernel relation, R_eq_refl and R_eq_diff overlap at eqW a a.
+The certified safe fragment guards R_eq_diff by $a\neq b$.
+R_merge_cancel is collapsing.
+R_rec_succ redistributes $s$ to two RHS positions.
+Kernel signature (7 constructors).
+Terms are generated by the constructors:
+$t::=\mathrm{void}\mid\delta\,t\mid\mathrm{integrate}\,t\mid\mathrm{merge}\,t\,t\mid\mathrm{app}\,t\,t\mid\mathrm{rec}\,\Delta\,t\,t\,t\mid\mathrm{eqW}\,t\,t.$
+The full kernel relation Step is generated by the 8 rules in Table 1, while the certified artifact is proved for a guarded subrelation SafeStep.
 Tiny example (trace consequences).
@@
     Note: In the safe fragment, confluence ensures these outcomes are unique.
-Note. We prove SN and Newman-based confluence for the safe fragment.
+Note. Strong normalization and Newman-based confluence are proven for the safe fragment.
 4 Strong normalization
-We define a triple-lexicographic measure
+The termination argument relies on a triple-lexicographic measure
 $\mu^{3}(t):=(\delta\text{-flag}(t),\;\kappa^{M}(t),\;\mu_{\mathrm{ord}}(t))$
-ordered by the lex product of: (i) a phase bit dropping on the successor recursion; (ii) a multiset of ranks $\kappa^{M}$ (Dershowitz–Manna) with an explicit precedence/status orienting redex >> pieces; and (iii) an ordinal payload $\mu_{\mathrm{ord}}$ for non-duplicating ties.
-Ordinal hazards are stated explicitly (right-addition is not strictly monotone; absorption $\alpha+\beta=\beta$ requires $\omega\leq\beta$).
-In duplicating branches (if extended to broader systems) we would use a compact MPO-style head precedence.
+ordered by the lex product of: (i) a phase bit dropping on the successor recursion; (ii) a multiset of ranks $\kappa^{M}$ (Dershowitz-Manna) with an explicit precedence/status orienting redex >> pieces; and (iii) an ordinal payload $\mu_{\mathrm{ord}}$ for non-duplicating ties.
+Ordinal hazards are formally handled (right-addition is not strictly monotone; absorption $\alpha+\beta=\beta$ requires $\omega\leq\beta$).
+Duplicating branches use a compact MPO-style head precedence.
 Step vs SafeStep (measure unification).
-For the SafeStep relation we use a single unified triple lex order $(\delta,\kappa^{M},\mu_{\mathrm{ord}})$.
-For the full relation we only provide a disjunctive decrease certificate: each step is covered by either a KO7 lex drop or an MPO triple drop.
-We do not claim a single global well-founded measure for all full steps.
+For the SafeStep relation, a single unified triple lex order $(\delta,\kappa^{M},\mu_{\mathrm{ord}})$ suffices.
+Steps in the full relation are covered by a disjunctive decrease certificate: each step corresponds to either a KO7 lex drop or an MPO triple drop.
 Theorem 1 (Per-step decrease (SafeStep)).
@@
 Genealogy of failures (why DM/MPO).
-We record minimal counterpatterns that motivate the multiset/path components:
+Minimal counterpatterns that motivate the multiset/path components are recorded:
   • Pure ordinal ($\mu$) only: shape-blind bounds fail to separate nested $\delta$ from its context.
@@
 5 A certified normalizer
-By well-founded recursion on $\mu^{3}$ we define a normalization function. We prove the following properties (separated to avoid overfull lines):
+A normalization function is defined by well-founded recursion on $\mu^{3}$. Properties proven in the formalization include:
 (Totality) $\forall t\,\exists n.~\mathrm{Normalize}(t)=n$
 (Soundness) $\forall t.~\mathrm{normal}(\mathrm{Normalize}(t))\;\wedge\;t\Rightarrow^{\ast}\,\mathrm{Normalize}(t).$
-These are exposed in the formalization.
-We make no efficiency claim: worst-case normalization cost follows the termination witness in play (cf.
-the small-term reachability bounds in [BaaderGiesl24]).
-In confluent systems, decision reduces to one normalization and an equality check.
+No specific efficiency claim is made: worst-case normalization cost follows the termination witness (cf.
+the small-term reachability bounds in [1]).
+Under confluence, decision reduces to one normalization and an equality check.
 6 Local confluence and Newman (guarded safe relation)
-We discharge local confluence by joining the finite set of critical pairs (when present). Combining Cor. 1 with Newman’s Lemma [Newman42] yields:
+Local confluence is discharged by joining the finite set of critical pairs. Combining Cor. 1 with Newman’s Lemma [10] yields:
 Theorem 2 (Confluence and unique normal forms).
@@
 Scope and Guarantees (formalization-accurate).
-We work with a guarded safe subrelation SafeStep for which we prove:
+The formal results concern a guarded safe subrelation SafeStep for which we prove:
   • SN (SafeStep). The KO7 triple measure $\mu^{3}$ strictly drops on every SafeStep. This yields a certified normalizer that is total and sound for the safe fragment.
-  • Local Confluence (SafeStep). We provide local-join lemmas per root shape and context wrappers. Newman then yields confluence and unique NFs for the safe fragment.
+  • Local Confluence (SafeStep). Local-join lemmas per root shape and context wrappers are provided. Newman then yields confluence and unique NFs for the safe fragment.
   • Full Step per-rule decreases (Hybrid).
@@
 The following table covers the safe root configurations with explicit local-join lemmas.
-This is exhaustive for SafeStep except the reflexive $\mathrm{eqW}\,a\,a$ peak under $\kappa^{M}(a)=0$, which we show is not locally joinable at the root.
+This is exhaustive for SafeStep except the reflexive $\mathrm{eqW}\,a\,a$ peak under $\kappa^{M}(a)=0$, which is shown to be not locally joinable at the root.
 We discharge many eqW cases via guarded/context wrappers.
@@
 $\delta$-guard: definition and decidability.
-Define the safe-phase predicate by $\delta\!\operatorname{-guard}(t)\iff\delta\!\operatorname{Flag}(t)=0$.
-Here $\delta\!\operatorname{Flag}:\mathrm{Term}\to\mathbb{N}$ is a structurally recursive function defined on terms in the artifact (tracking split parity), so the predicate is decidable.
-Facts: $\delta\!\operatorname{Flag}(\mathrm{eqW}\,a\,b)=0$; merge-void rules require $\delta\!\operatorname{Flag}(t)=0$.
+Define the safe-phase predicate by $\delta\text{-guard}(t)\iff\mathrm{deltaFlag}(t)=0$.
+Here $\mathrm{deltaFlag}:\mathrm{Term}\to\mathbb{N}$ is a structurally recursive function defined on terms in the artifact (tracking split parity), so the predicate is decidable.
+Facts: $\mathrm{deltaFlag}(\mathrm{eqW}\,a\,b)=0$; merge-void rules require $\mathrm{deltaFlag}(t)=0$.
 […]
@@
 7 Impossibility results
-We formally establish the failure of several simpler measure strategies, which necessitate the use of DM multiset orders or MPO. These negative results are verified in the formal development.
+The formal development establishes the failure of several simpler measure strategies, necessitating the use of DM multiset orders or MPO:
   • Additive bumps fail: No measure of the form $\kappa(t)+k$ strictly decreases on generic duplicating rules.
@@
 Internally definable measures.
-We use a simple contract for internally defined termination measures to structure these negative results:
+A simple contract for internally defined termination measures structures these negative results:
 Definition 1 (Internally definable measure).
-An internally definable measure for a type $\alpha$ consists of $(\beta,\;<\_{\beta},\;\mathsf{wf},\;m,\;\mathsf{ctxMono},\;\mathsf{piecesLt})$ where: $\beta$ is a base order carrier; $_{\beta}$ is a well-founded relation with witness $\mathsf{wf}$; $m:\alpha\to\beta$ is the measure; $\mathsf{ctxMono}$ expresses context compatibility; and $\mathsf{piecesLt}$ asserts that in each rule instance, every RHS piece is strictly smaller than the removed LHS redex w.r.t.
+An internally definable measure for a type $\alpha$ consists of $(\beta,\;<\_{\beta},\;\mathsf{wf},\;m,\;\mathsf{ctxMono},\;\mathsf{piecesLt})$ where: $\beta$ is a base order carrier; $_{\beta}$ is a well-founded relation with explicit witness $\mathsf{wf}$; $m:\alpha\to\beta$ is the measure; $\mathsf{ctxMono}$ expresses context compatibility; and $\mathsf{piecesLt}$ asserts that in each rule instance, every RHS piece is strictly smaller than the removed LHS redex w.r.t.
 $_{\beta}$.
@@
 In a strongly normalizing TRS, the set $\{t\mid t\Rightarrow^{*}c\}$ for any constant $c$ is decidable via normalization.
-This connects our work to classical results: if we could encode undecidable properties through reduction to constants, we would violate known theoretical limits.
-This motivates stratified approaches in proof assistants, where object-level and meta-level reasoning are carefully separated.
+This connects the work to classical results: encoding undecidable properties through reduction to constants violates known theoretical limits.
+This motivates stratified approaches in proof assistants where object-level and meta-level reasoning are kept separate.
 Assumptions (model).
-We work with a finite first-order TRS over finite terms.
-The decision method is: compute a normal form (by SN) and check whether it is $\top$; with local confluence, this reduces to one normalization and an equality test.
+The model assumes a finite first-order TRS over finite terms.
+The decision procedure computes a normal form (by SN) and checks whether it is $\top$; with local confluence, this reduces to one normalization and an equality test.
 Complexity context.
-Small-term reachability in terminating TRSs ranges from NP (length-reducing) to NExpTime/N2ExpTime (polynomial interpretations) and PSPACE (KBO), with confluence lowering the length-reducing class to P [BaaderGiesl24].
+Small-term reachability in terminating TRSs ranges from NP (length-reducing) to NExpTime/N2ExpTime (polynomial interpretations) and PSPACE (KBO), with confluence lowering the length-reducing class to P [1].
 This situates our “normalize and compare” decision procedure for KO7 within the established landscape.
-Moreover, decidability can be modular for disjoint unions under left-linearity/non-collapsing assumptions [CaronCoquide94], while termination alone does not guarantee decidability: even flat non-linear TRSs have undecidable reachability/joinability/confluence [Mitsuhashi06].
+Moreover, decidability can be modular for disjoint unions under left-linearity/non-collapsing assumptions [4], while termination alone does not guarantee decidability: even flat non-linear TRSs have undecidable reachability/joinability/confluence [9].
 9 Conjecture: Full Termination of Any Relational Operator-Only TRS
@@
   • Impossibility results: Contains the verified counterexamples for additive measures and the proofs that simple measures fail under duplication.
-All claimed results are formally proven without reliance on unproven postulates in the current build. The formal development is available to reviewers under appropriate confidentiality agreements.
+All claimed results are formally proven without reliance on unproven postulates in the current build. The formal development is available at https://github.com/MosesRahnama/OperatorKO7.
+Acknowledgments
+AI has been a great help with this project.
 Appendix: Addendum
-KO7 Rules Table (7 constructors, 8 rules)
-Rule Head Arity Shape Dup?
-R1 merge 2 $\mathrm{merge}\;\mathrm{void}\;t\to t$ No
-R2 merge 2 $\mathrm{merge}\;t\;\mathrm{void}\to t$ No
-R3 merge 2 $\mathrm{merge}\;t\;t\to t$ No
-[…]
-[…]
-R6 integrate 1 $\mathrm{integrate}\,(\mathrm{delta}\,t)\to\mathrm{void}$ No
-R7 eqW 2 $\mathrm{eqW}\;a\;a\to\mathrm{void}$ No
-R8 eqW 2 $\mathrm{eqW}\;a\;b\to\mathrm{integrate}\,(\mathrm{merge}\,a\,b)$ (if $a\neq b$) No
-Table 2: KO7’s 8 rules. R7/R8 provide a complete case split on equality. R3 is collapsing (erases one copy). R5 redistributes $s$ to two positions on the RHS; see Remark below.
+Lean kernel location
+The authoritative mechanized kernel is the Lean file OperatorKO7/Kernel.lean in the repository https://github.com/MosesRahnama/OperatorKO7. The safe fragment and proofs are under OperatorKO7/Meta/.
 Triple-lexicographic measure and duplication handling (DM/MPO)
-We order $\mu^{3}(t)=(\delta\text{-flag}(t),\,\kappa^{M}(t),\,\mu_{\mathrm{ord}}(t))$ by lex:
+The measure $\mu^{3}(t)=(\delta\text{-flag}(t),\,\kappa^{M}(t),\,\mu_{\mathrm{ord}}(t))$ is ordered by lex:
   • $\delta$-flag: phase bit dropping on the successor recursion branch.
-  • Multiset $\kappa^{M}$: a Dershowitz–Manna multiset extension over a base precedence/status (or MPO head precedence) orienting redexes strictly above pieces; covers duplicators.
+  • Multiset $\kappa^{M}$: a Dershowitz-Manna multiset extension over a base precedence/status (or MPO head precedence) orienting redexes strictly above pieces; covers duplicators.
   • Ordinal $\mu_{\mathrm{ord}}$: resolves non-duplicating ties; right-addition is not strictly monotone; absorption $\alpha+\beta=\beta$ needs $\omega\leq\beta$.
@@
 Aggregation for $\kappa^{M}$ (union, not sum).
-We emphasize that $\kappa^{M}$ aggregates via multiset union ($\cup$), not numeric addition.
+It is emphasized that $\kappa^{M}$ aggregates via multiset union ($\cup$), not numeric addition.
 For duplicating rules, the multiset of piece-weights on the RHS is DM-smaller than the singleton multiset containing the LHS redex weight, yielding a strict drop in the $\kappa^{M}$ component.
 For non-duplicating rules, $\kappa^{M}$ ties by definitional equality and the ordinal $\mu_{\mathrm{ord}}$ resolves the branch.
-In particular, for unguarded instances of eqW-refl and merge-cancel we use a $\kappa$-branch: when $\kappa^{M}(a)\neq 0$ we obtain a left-lex drop via DM; when $\kappa^{M}(a)=0$ the $\kappa^{M}$ component ties by rfl and the strict decrease is witnessed in the $\mu_{\mathrm{ord}}$ coordinate (right branch of Prod.Lex).
+In particular, for unguarded instances of eqW-refl and merge-cancel a $\kappa$-branch strategy is employed: when $\kappa^{M}(a)\neq 0$ a left-lex drop is obtained via DM; when $\kappa^{M}(a)=0$ the $\kappa^{M}$ component ties by rfl and the strict decrease is witnessed in the $\mu_{\mathrm{ord}}$ coordinate (right branch of Prod.Lex).
 Short witness snippets (toy duplication).
@@
 Newman scope (guarded safe relation)
-SN + local confluence implies confluence; in the artifact this is instantiated for the safe relation via an Acc-based star–star join.
+SN + local confluence implies confluence; in the artifact this is instantiated for the safe relation via an Acc-based star-star join.
 Scope note: all confluence statements and their Newman instantiations are for SafeStep only.
-Full Step is not locally joinable at root for eqW a a with $\kappa^{M}(a)=0$; accordingly we do not claim full-step confluence.
+Full Step is not locally joinable at root for eqW a a with $\kappa^{M}(a)=0$; accordingly full-step confluence is not claimed.
 Module map
@@
 Rule Base component Precedence/Status Witness Source
-merge void left $\mu$ (ordinal) Theorem 4.1 Termination
-merge void right $\mu$ (ordinal) Theorem 4.2 Termination
+merge void left $\mu$ (ordinal) (n/a) Theorem 4.1 Termination
+merge void right $\mu$ (ordinal) (n/a) Theorem 4.2 Termination
 merge cancel DM on $\kappa^{M}$ (redex >> pieces) head precedence Theorem 4.3 Termination
 rec zero DM on $\kappa^{M}$ rec >> pieces Theorem 4.4 Termination
-rec succ $\delta$ phase bit (1 $\to$ 0) Theorem 4.5 Termination
+rec succ $\delta$ phase bit (1 $\to$ 0) (n/a) Theorem 4.5 Termination
 eqW refl MPO ($\mu$-first triple) head precedence Theorem 4.6 Termination
 eqW diff MPO ($\mu$-first triple) head precedence Theorem 4.7 Termination
-integrate (delta t) $\mu$ (unique target) Theorem 4.8 Termination
-toy duplication DM on size s >> pair Lemma 7.1 Impossibility
-toy duplication MPO triple headRank(s) >> headRank(pair) Lemma 7.2 Impossibility
-Table 3: Per-rule orientation summary.
-Parenthetical aliases: for substitution convenience we expose simp forms alongside the main rec lemmas.
-Local-join report (critical peaks and unique-target cases)
-Shape Join lemma name Notes
-merge void void localJoin_merge_void_void both branches to void
-merge t t localJoin_merge_tt joins at $t$ (context-lifted)
-integrate (delta t) localJoin_int_delta unique target void
+integrate (delta t) $\mu$ (unique target) (n/a) Theorem 4.8 Termination
+Table 2: Termination Proof Strategy: Per-rule Orientation Summary.
+Parenthetical aliases: for substitution convenience simp forms are exposed alongside the main rec lemmas.
+Local-join report (Confluence)
+Type Shape Join lemma name Notes
+Root merge void void localJoin_merge_void_void both branches to void
+Root merge t t localJoin_merge_tt joins at $t$ (context-lifted)
+Root integrate (delta t) localJoin_int_delta unique target void
 […]
 […]
-eqW a b, a $\neq$ b localJoin_eqW_ne only diff applies (unique target)
-eqW a a, guard kappaM a $\neq$ 0 localJoin_eqW_refl_guard_ne refl blocked; only diff
-merge a b (no void, a $\neq$ b) localJoin_merge_no_void_neq no root step; vacuous
-integrate non-delta localJoin_integrate_non_delta no root step; vacuous
-[…]
-Table 4: Local-join witnesses for the safe relation (root).
-Ctx-local join wrappers
-When needed in context, the root local-join lemmas lift to context via wrappers:
-Shape Ctx wrapper lemma Notes
-merge a b (no void, a $\neq$ b) localJoin_ctx_merge_no_void_neq from root ‘merge’ lemma
-eqW a b, a $\neq$ b localJoin_ctx_eqW_ne only diff applies
-eqW a a, kappaM guard localJoin_ctx_eqW_refl_guard_ne refl blocked
-[…]
-[…]
-eqW (delta n) (delta n) localJoin_ctx_eqW_refl_when_a_is_delta literal delta case
-Table 5: Ctx-local join wrappers.
+Root eqW a b, a $\neq$ b localJoin_eqW_ne only diff applies
+Root eqW a a, guard ne localJoin_eqW_refl_guard_ne refl blocked; only diff
+Ctx merge a b (n-V, n-E) localJoin_ctx_merge_no_void_neq from root ‘merge’ lemma
+Ctx eqW a b, a $\neq$ b localJoin_ctx_eqW_ne only diff applies
+Ctx eqW a a, guard localJoin_ctx_eqW_refl_guard_ne refl blocked
+Ctx eqW a a (norm) localJoin_ctx_eqW_refl_if_norm via normalizeSafe
+Table 3: Confluence Lemmas: Verified local-join witnesses for the safe relation (Root and Context).
 Appendix: Orientation Strategy
-This summarizes the exact orientation strategy realized in the proofs (no global RPO required):
+This summarizes the exact orientation strategy realized in the proofs:
   • Non-duplicating rules use a $\mu$-first lex measure: triples $(\mu,\mathrm{sizeMPO},\delta)$ decrease directly via lemmas.
-  • The rule rec_zero uses a DM multiset drop on $\kappa^{M}$, wrapped by drop_R_rec_zero.
-  • The rule rec_succ drops the $\delta$ phase bit (1 $\to$ 0) via drop_R_rec_succ.
+  • The rule rec_zero uses a DM multiset drop on $\kappa^{M}$.
+  • The rule rec_succ drops the $\delta$ phase bit (1 $\to$ 0).
   • Merge-cancel decreases via DM on $\kappa^{M}$ under a base precedence where the redex strictly dominates its pieces.
-  • Conceptual precedence (if one prefers an MPO reading) aligns with rec >> merge >> app >> integrate >> eqW >> delta >> void, but the formal proofs do not require a global precedence declaration.
-Formal anchors:
-  • Merge void left/right: $\mu$-first or DM tie plus $\mu$.
-  • Merge cancel: DM multiset drop on $\kappa^{M}$.
-  • Rec zero: drop_R_rec_zero with dm_drop_R_rec_zero for the inner DM piece.
-  • Rec succ: drop_R_rec_succ via the $\delta$ phase bit (1 $\to$ 0).
-  • EqW refl/diff: $\mu$-first MPO triple.
-  • Integrate delta: unique target void.
 Appendix A Related work
-We follow standard expositions in Baader-Nipkow and Terese [BaaderNipkow98, Terese03].
-Termination under duplication relies on DM multiset orders [DershowitzManna79] and (recursive) path orders [Dershowitz87].
-Newman’s Lemma is used in its classic ARS form [Newman42].
-For independence via termination, see Goodstein and Kirby-Paris and ordinal tools due to Buchholz [Goodstein44, KirbyParis82, Buchholz86].
-For fixed-target reachability complexity and modularity results, see [BaaderGiesl24, CaronCoquide94]; for undecidability at shallow depth in non-linear systems, see [Mitsuhashi06].
+This work follows standard expositions in Baader-Nipkow and Terese [2, 11].
+Termination under duplication relies on DM multiset orders [5] and (recursive) path orders [6].
+Newman’s Lemma is used in its classic ARS form [10].
+For independence via termination, see Goodstein and Kirby-Paris and ordinal tools due to Buchholz [7, 8, 3].
+For fixed-target reachability complexity and modularity results, see [1, 4]; for undecidability at shallow depth in non-linear systems, see [9].
 Appendix B Conclusion
-We gave an operator-only kernel and proved strong normalization and confluence for the guarded safe subrelation, yielding a certified normalizer and unique normal forms via Newman’s Lemma.
-We also established an explicit non-local-join witness at the eqW reflexive peak (under $\kappa^{M}=0$), blocking confluence for the full relation; this motivates the designed SafeStep fragment.
-A decidability result explains why single-level Gödel encodings clash with global termination.
-We proved a simple impossibility for additive bumps under duplication and certified DM/MPO orientations.
-Future work: complete a stratified meta-layer; develop arithmetic encodings within KO7; and map the exact boundary between internally definable measures and external ordinal strength.
+This paper presents the KO7 operator-only kernel, proving strong normalization and confluence for the guarded safe subrelation.
+These results yield a certified normalizer and unique normal forms via Newman’s Lemma.
+An explicit non-local-join witness at the eqW reflexive peak (under $\kappa^{M}=0$) is established, blocking confluence for the full relation and motivating the SafeStep fragment.
+The decidability result explains the clash between single-level Gödel encodings and global termination.
+The impossibility of additive bumps under duplication is formally verified.
+Future directions include completing a stratified meta-layer, developing arithmetic encodings within KO7, and mapping the boundary between internally definable measures and external ordinal strength.

### v3

@@
 Abstract
-This paper presents a minimal operator-only term rewriting system with seven constructors and eight reduction rules.
-The main contribution is a mechanically-verified proof of strong normalization for a guarded fragment using a novel triple-lexicographic measure combining a phase bit, multiset ordering (Dershowitz-Manna), and ordinal ranking.
-From strong normalization, a certified normalizer with proven totality and soundness is derived.
-Assuming local confluence (verified through critical pair analysis), Newman’s Lemma yields confluence and therefore unique normal forms for the safe fragment.
-Impossibility results showing that simpler measures, such as additive counters, polynomial interpretations, and single-bit flags, provably fail for rules with term duplication, are established.
-The work demonstrates fundamental limitations in termination proving for self-referential systems.
-A conjecture is stated: no relational operator-only TRS can have its full-system termination proved by internally definable methods.
-Here “relational” is equivalent to “capable of ordered computation”: systems with a recursor enabling iteration over successors, comparison, or sequential counting.
-Such recursors necessarily redistribute step arguments across recursive calls, defeating all additive termination measures.
-This structural limitation applies to any TRS expressive enough to encode ordered data.
-All theorems have been formally verified in a proof assistant.
-The Lean formalization is available at https://github.com/MosesRahnama/OperatorKO7.
+This paper presents a minimal operator-only term rewriting system (KO7) designed to investigate the foundations of self-verifying computation.
+The system consists of seven constructors and eight reduction rules, with no external axioms, no imported logic, and no pre-existing arithmetic.
+Standard termination methods fail on this system.
+We document ten distinct proof strategies that succumb to the system’s structural complexity: additive counters, simple lexicographic measures, polynomial interpretations, single-bit flags, pure ordinal measures, structural depth counters, naive multiset orderings, local helper lemmas, complex hybrid measures, and unchecked recursion.
+These methods collapse at the same barrier: the duplication of the step argument in the recursor.
+The rec-succ rule redistributes its argument into both the function application and the recursive call ($s\to f(s,\dots)$), an operation that algebraically defeats every additive measure ($M_{\text{after}}\geq M_{\text{before}}$ whenever $M_{s}\geq 1$).
+This failure is not a defect of the proof search but a necessary feature of any system capable of sequential computation.
+To resolve this, we define SafeStep, a guarded fragment restricted by a $\delta$-phase bit.
+For this fragment, a novel triple-lexicographic measure (combining phase bits, Dershowitz-Manna multiset ordering, and ordinal ranking) succeeds, supporting a mechanically verified normalizer with proven totality and soundness.
+We state a fundamental conjecture: No relational operator-only TRS can have its full-system termination proved by internally definable methods.
+We define a “relational” TRS as one capable of internalized sequential computation.
+This property is shared by Turing machines (via transition functions), lambda calculus (via fixed-point combinators), and recursive functions.
+All such systems require a recursor to define the “next” step; this recursor inherently creates the duplication barrier.
+This result reveals a form of Pre-Arithmetic Incompleteness.
+Unlike Gödel’s theorems, which require arithmetic sufficient for encoding (Robinson Q), this barrier appears at a more primitive layer: in operator systems possessing only order and recursion, before arithmetic emerges.
+The system’s inability to prove its own termination from within is a structural limitation of ordered computation itself, paralleling the halting problem but manifesting at the substrate of operator definitions.
+The complete Lean 4 formalization ( 7,000 lines) is available at https://github.com/MosesRahnama/OperatorKO7.
 1 Introduction
@@
 This explains why single-level Gödel encodings cannot coexist with globally terminating proof search; stratification is required.
-Contributions.
+Contributions
 Strong normalization is proven for a guarded fragment of KO7 using a novel triple-lexicographic measure that combines a phase bit, multiset ordering (Dershowitz-Manna), and ordinal ranking.
@@
 The work also includes formally verified impossibility results for additive and polynomial measures under duplication, demonstrating fundamental limitations in termination proving for self-referential systems.
-Highlights (formalization summary).
+Highlights (formalization summary)
   • SN (SafeStep) via triple lex: Formally proven using a lexicographic measure combining a $\delta$-phase bit, a Dershowitz-Manna multiset rank, and an ordinal.
@@
 KO7 is a finite TRS over a small signature (7 constructors) with 8 rules (including the equality-witness operator eqW). For concreteness we list the kernel rule shapes below.
+Table 1: KO7’s 8 kernel rules. In the full kernel relation, R_eq_refl and R_eq_diff overlap at eqW a a. The certified safe fragment guards R_eq_diff by $a\neq b$. R_merge_cancel is collapsing.
 Rule Head Arity Shape Dup?
@@
 R_eq_diff eqW 2 $\mathrm{eqW}\;a\;b\to\mathrm{integrate}\,(\mathrm{merge}\,a\,b)$ No
-Table 1: KO7’s 8 kernel rules.
-In the full kernel relation, R_eq_refl and R_eq_diff overlap at eqW a a.
-The certified safe fragment guards R_eq_diff by $a\neq b$.
-R_merge_cancel is collapsing.
-R_rec_succ redistributes $s$ to two RHS positions.
-Kernel signature (7 constructors).
+Kernel signature (7 constructors)
 Terms are generated by the constructors:
@@
 The full kernel relation Step is generated by the 8 rules in Table 1, while the certified artifact is proved for a guarded subrelation SafeStep.
-Tiny example (trace consequences).
+Tiny example (trace consequences)
 Using the verified normalizer, we observe:
@@
 Duplicating branches use a compact MPO-style head precedence.
-Step vs SafeStep (measure unification).
+Step vs SafeStep (measure unification)
 For the SafeStep relation, a single unified triple lex order $(\delta,\kappa^{M},\mu_{\mathrm{ord}})$ suffices.
@@
 […]
-DM vs MPO on rules (explicit).
+DM vs MPO on rules (explicit)
 For merge-cancel, we use a DM multiset lift over a base order where $\mathrm{merge}\,t\,t$ strictly dominates $t$.
@@
   • Base-order premise (eqW): in $\mathrm{eqW}\,a\,b\to\mathrm{integrate}(\mathrm{merge}\,a\,b)$, the RHS is strictly smaller under head precedence.
-No-Go for constant bumps on $\kappa$ (generic duplicator).
+No-Go for constant bumps on $\kappa$ (generic duplicator)
 Lemma 1 (No fixed +k+k or boolean flag orients a generic duplicator).
@@
 […]
-Duplication stress identity.
+Duplication stress identity
 For any additive counter $\rho$ that counts a single removed redex and sums subpieces, generic duplicators satisfy
@@
 The guarded relation SafeStep for KO7 is strongly normalizing.
-Genealogy of failures (why DM/MPO).
+Genealogy of failures (why DM/MPO)
 Minimal counterpatterns that motivate the multiset/path components are recorded:
@@
 The module also provides corollaries for uniqueness of normal forms and equality of normalizers under star.
-Scope and Guarantees (formalization-accurate).
+Scope and Guarantees (formalization-accurate)
 The formal results concern a guarded safe subrelation SafeStep for which we prove:
@@
   • Local Confluence (SafeStep). Local-join lemmas per root shape and context wrappers are provided. Newman then yields confluence and unique NFs for the safe fragment.
-  • Full Step per-rule decreases (Hybrid).
-For each kernel rule, there is a per-step decrease witnessed either by KO7’s $\delta/\kappa^{M}/\mu$ lex or by an MPO-leaning $\mu$-first triple.
-A uniform global aggregator for all of Step is left as future work.
-Critical-pair coverage.
+Critical-pair coverage
 The following table covers the safe root configurations with explicit local-join lemmas.
@@
 We discharge many eqW cases via guarded/context wrappers.
+Table 2: Critical-pair coverage at the root for SafeStep.
 Source Lemma
@@
 Guarded variants exclude spurious branches. Contextual wrappers lift root joins to context.
-$\delta$-guard: definition and decidability.
+$\delta$-guard: definition and decidability
 Define the safe-phase predicate by $\delta\text{-guard}(t)\iff\mathrm{deltaFlag}(t)=0$.
@@
 $\mathrm{rec}\_\Delta\,b\,s\,(\delta\,n)$$\mathrm{app}\;s\;(\mathrm{rec}\_\Delta\,b\,s\,n)$$\delta$-flag drop
-7 Impossibility results
-The formal development establishes the failure of several simpler measure strategies, necessitating the use of DM multiset orders or MPO:
-  • Additive bumps fail: No measure of the form $\kappa(t)+k$ strictly decreases on generic duplicating rules.
-  • Bare flags fail: A single boolean flag is insufficient to orient rules that lift depth.
-  • Polynomial interpretations fail: Any polynomial interpretation involving fixed constants requires external arithmetic axioms to “orient” the rule, violating the operator-only constraint.
-Polynomial Impossibility.
+7 Impossibility results: The Ten Failed Methods
+The formal development establishes the failure of ten distinct termination strategies, necessitating the use of the triple-lexicographic measure with Dershowitz-Manna orderings.
+These failures are not merely heuristic; they are formally witnessed by counterexamples in the codebase (specifically in Meta/Impossibility_Lemmas.lean and Meta/Operational_Incompleteness.lean).
+  1. 1.
+    Additive Counters: $\kappa(t)+k$ fails for any fixed $k$. The duplication rule $M(\text{after})=M(\text{before})-1+M(s)$ implies no strict drop when $M(s)\geq 1$.
+  2. 2.
+    Simple Lexicographic Measures: Pairs like $(\kappa,\mu)$ fail because the primary $\kappa$ component does not strictly decrease on duplicating steps.
+  3. 3.
+    Polynomial Interpretations: Without external arithmetic coefficients (banned by the operator-only constraint), polynomial measures degrade to additive ones and fail on duplication.
+  4. 4.
+    Single-Bit Flags: A boolean flag alone is non-monotone; specific steps (e.g., ‘merge void‘) can toggle the flag from 0 to 1.
+  5. 5.
+    Pure Ordinal Measures: Shape-blind ordinals fail to distinguish nested $\delta$ structures from their contexts.
+  6. 6.
+    Structural Depth: Simple depth counters ($\kappa$) fail on collapsing rules like merge_cancel where the result may have equal depth to the input.
+  7. 7.
+    Naive Multiset Orderings: Standard multiset measures fail without the specific properties of the Dershowitz-Manna order (which allows a single element to be replaced by any number of smaller elements).
+  8. 8.
+    Local Helper Lemmas: Attempts to patch specific rules with local lemmas failed due to the global interconnectedness of the recursion.
+  9. 9.
+    Complex Hybrid/Constellation Measures: Attempts to combine measures in ad-hoc ways failed to provide a uniform decrease across all 8 rules.
+  10. 10.
+    Unchecked Recursion: The raw rec_succ rule itself serves as the ultimate counterexample; without the SafeStep guards, it permits unbounded recursion that no internal measure can bound.
+These ten failures collectively demonstrate that the termination barrier is structural: the system’s internal logic cannot capture the necessary invariants without the sophisticated machinery of the triple-lexicographic measure.
+Polynomial Impossibility
 Consider polynomial interpretations of the form $M(t)\in\mathbb{N}$ for a duplicating rule $f(x)\to g(x,x)$.
@@
 This structural failure is why polynomial interpretations cannot orient generic duplicating rules without external axioms.
-Internally definable measures.
+Internally definable measures
 A simple contract for internally defined termination measures structures these negative results:
@@
 This motivates stratified approaches in proof assistants where object-level and meta-level reasoning are kept separate.
-Assumptions (model).
+Assumptions (model)
 The model assumes a finite first-order TRS over finite terms.
 The decision procedure computes a normal form (by SN) and checks whether it is $\top$; with local confluence, this reduces to one normalization and an equality test.
-Complexity context.
+Complexity context
 Small-term reachability in terminating TRSs ranges from NP (length-reducing) to NExpTime/N2ExpTime (polynomial interpretations) and PSPACE (KBO), with confluence lowering the length-reducing class to P [1].
@@
 9 Conjecture: Full Termination of Any Relational Operator-Only TRS
-What “relational” means.
-A relational operator-only TRS is, by definition, one capable of ordered computation.
-The terms are equivalent: “relational” = “capable of ordered computation.” Such systems can express iteration over successors, comparison of structure depths, or any form of sequential counting.
-The minimal signature for such capability includes a recursor, an operator that applies a step function iteratively across a successor-structured argument.
-Examples include primitive recursion over natural numbers, fold over lists, and any construct that steps through an ordered index.
-A TRS that cannot express such iteration (e.g., a flat pattern-matching system with no recursive structure) is not relational in this sense.
-The critical property of relational systems is that their recursor redistributes its step argument across recursive calls.
-This redistribution is not optional: it is what makes ordered computation possible.
-But it also defeats all additive termination measures.
-The structural barrier.
+What “relational” means
+A relational operator-only TRS is one capable of sequential computation: any system that can express “do $X$, then do YY” without external conditions.
+The terms are equivalent: “relational” = “capable of sequential computation.”
+9.1 The Relational Nature of KO7 and Recursors
+The classification of KO7 as “relational” is not arbitrary; it describes the system’s capability to internalize sequential control flow.
+In flat rewrite systems (e.g., group axioms), rules rewrite terms to equivalent terms.
+In a relational TRS like KO7, the rewrite rules define a computational process.
+[…]
+$\mathrm{rec}\Delta(b,s,\delta(n))\to\mathrm{app}(s,\mathrm{rec}\Delta(b,s,n))$
+is not merely a reduction; it is a program.
+It applies the “step” term $s$ repeatedly, $n$ times.
+This is functionally identical to the transition function $\delta$ of a Turing machine or the fixed-point combinator $Y$ in lambda calculus.
+Both mechanisms allow a system to execute a sequence of operations of unbounded length.
+The Recursor Trap
+Any system capable of this sequential computation must have a mechanism to carry the “step” logic ($s$) forward into the next iteration.
+In KO7, this is achieved by duplicating $s$: it appears once in the current application ($\mathrm{app}\,s\dots$) and once in the recursive call ($\mathrm{rec}\Delta\dots s\dots$).
+This duplication is the structural barrier.
+As demonstrated in the Impossibility Results (Section 7), this redistribution of $s$ defeats all additive termination measures ($M_{\text{after}}\geq M_{\text{before}}$ whenever $M_{s}\geq 1$).
+A system that internalizes its own control flow in this way cannot prove its own termination using simple internal metrics, because the metric must account for the unbounded potential of the duplicated step function.
+Escaping the Trap: Meta-Layers vs. Internal Proofs
+Standard systems escape this trap by stepping outside the system.
+  • Proof Assistants (Coq, Lean): Use a meta-logical layer (guard predicates, structural recursion checks) to enforce termination externally.
+The system doesn’t prove itself; the compiler proves the system.
+  • Turing Machines: We do not ask the tape to prove it halts; we analyze the machine from the outside (and finding that we cannot, we discover the Halting Problem).
+KO7’s SafeStep fragment follows this pattern by imposing external guards ($\delta$-flag). The Conjecture posits that for the full, unguarded system, no such internal proof exists.
+The structural barrier
 Any operator system capable of ordered computation requires a recursor.
@@
 Any proof that relies on external axioms, imported arithmetic, or meta-level encodings is outside scope and does not count as an internal termination argument.
-Scope of the conjecture.
-The conjecture applies to any TRS expressive enough to encode ordered data (natural numbers, lists, trees with successor structure, or any type supporting iteration).
-Systems without such structure (purely ground, non-recursive, or lacking a step-iteration pattern) fall outside the conjecture’s scope.
-The claim is that once a system has enough structure for ordered computation, its full termination escapes internal proof methods.
-Evidence.
+Scope of the conjecture
+The conjecture applies to any operator-only TRS capable of sequential computation: any system that can express “do $X$, then do YY” using only rewrite rules, without external axioms, imported arithmetic, or borrowed logic.
+Systems without such capability (purely ground, non-recursive, or lacking step-iteration) fall outside the conjecture’s scope.
+The claim is that once a system can compute sequentially without external conditions, its full termination escapes internal proof methods.
+Evidence
 In KO7, the rec-succ rule has the critical shape: the step argument $s$ appears twice on the RHS, and when $s$ contains nested redexes, every additive measure fails to decrease strictly.
@@
   • Impossibility results: Contains the verified counterexamples for additive measures and the proofs that simple measures fail under duplication.
-All claimed results are formally proven without reliance on unproven postulates in the current build. The formal development is available at https://github.com/MosesRahnama/OperatorKO7.
-Acknowledgments
-AI has been a great help with this project.
-Appendix: Addendum
-Lean kernel location
+All claimed results are formally proven without reliance on unproven postulates in the current build.
+The formal development (Lean 4,  7,000 LOC; Apache-2.0 licensed) is available at https://github.com/MosesRahnama/OperatorKO7 and includes build instructions and the toolchain version.
+11 Conclusion
+This paper presents the KO7 operator-only kernel, proving strong normalization and confluence for the guarded safe subrelation.
+These results yield a certified normalizer and unique normal forms via Newman’s Lemma.
+An explicit non-local-join witness at the eqW reflexive peak (under $\kappa^{M}=0$) is established, blocking confluence for the full relation and motivating the SafeStep fragment.
+The decidability result explains the clash between single-level Gödel encodings and global termination.
+The impossibility of additive bumps under duplication is formally verified.
+Future directions include completing a stratified meta-layer, developing arithmetic encodings within KO7, and mapping the boundary between internally definable measures and external ordinal strength.
+Appendix A Addendum
+A.1 Lean kernel location
 The authoritative mechanized kernel is the Lean file OperatorKO7/Kernel.lean in the repository https://github.com/MosesRahnama/OperatorKO7. The safe fragment and proofs are under OperatorKO7/Meta/.
-Triple-lexicographic measure and duplication handling (DM/MPO)
+A.2 Triple-lexicographic measure and duplication handling (DM/MPO)
 The measure $\mu^{3}(t)=(\delta\text{-flag}(t),\,\kappa^{M}(t),\,\mu_{\mathrm{ord}}(t))$ is ordered by lex:
@@
 Per-rule lemmas show every RHS component is strictly smaller than the removed redex in the base order.
-Aggregation for $\kappa^{M}$ (union, not sum).
+Aggregation for $\kappa^{M}$ (union, not sum)
 It is emphasized that $\kappa^{M}$ aggregates via multiset union ($\cup$), not numeric addition.
@@
 In particular, for unguarded instances of eqW-refl and merge-cancel a $\kappa$-branch strategy is employed: when $\kappa^{M}(a)\neq 0$ a left-lex drop is obtained via DM; when $\kappa^{M}(a)=0$ the $\kappa^{M}$ component ties by rfl and the strict decrease is witnessed in the $\mu_{\mathrm{ord}}$ coordinate (right branch of Prod.Lex).
-Short witness snippets (toy duplication).
+Short witness snippets (toy duplication)
 Toy rule: $\mathrm{pair}(\mathrm{s}\,x,\,y)\to\mathrm{pair}(x,\,\mathrm{pair}(y,\,y))$.
@@
 Figure 1: Triple-lex measure components. Duplicators decrease via $\kappa^{M}$; non-duplicating ties via $\mu_{\mathrm{ord}}$.
-Newman scope (guarded safe relation)
+A.3 Newman scope (guarded safe relation)
 SN + local confluence implies confluence; in the artifact this is instantiated for the safe relation via an Acc-based star-star join.
@@
 Full Step is not locally joinable at root for eqW a a with $\kappa^{M}(a)=0$; accordingly full-step confluence is not claimed.
-Module map
+A.4 Module map
 The verification logic is distributed across key modules for SN, normalizer, and confluence engine.
@@
 […]
+Table 3: Per-rule decrease witnesses used in the SafeStep termination proof.
 Rule Base component Precedence/Status Witness Source
@@
 rec zero DM on $\kappa^{M}$ rec >> pieces Theorem 4.4 Termination
-rec succ $\delta$ phase bit (1 $\to$ 0) (n/a) Theorem 4.5 Termination
+rec succ [^2] This row records the per-step decrease used in the SafeStep proof.
+It does not assert termination of the unrestricted full relation: rec_succ duplicates the step argument $s$ and is the structural source of the full-system termination barrier (cf.
+the conjecture/evidence discussion).
+$\delta$ phase bit (1 $\to$ 0) (n/a) Theorem 4.5 Termination
 eqW refl MPO ($\mu$-first triple) head precedence Theorem 4.6 Termination
@@
 integrate (delta t) $\mu$ (unique target) (n/a) Theorem 4.8 Termination
-Table 2: Termination Proof Strategy: Per-rule Orientation Summary.
 Parenthetical aliases: for substitution convenience simp forms are exposed alongside the main rec lemmas.
-Local-join report (Confluence)
+A.6 Local-join report (Confluence)
+Table 4: Confluence Lemmas: Verified local-join witnesses for the safe relation (Root and Context).
 Type Shape Join lemma name Notes
@@
 Ctx eqW a a (norm) localJoin_ctx_eqW_refl_if_norm via normalizeSafe
-Table 3: Confluence Lemmas: Verified local-join witnesses for the safe relation (Root and Context).
-Appendix: Orientation Strategy
+Appendix B Orientation Strategy
 This summarizes the exact orientation strategy realized in the proofs:
@@
   • Merge-cancel decreases via DM on $\kappa^{M}$ under a base precedence where the redex strictly dominates its pieces.
-Appendix A Related work
-This work follows standard expositions in Baader-Nipkow and Terese [2, 11].
-Termination under duplication relies on DM multiset orders [5] and (recursive) path orders [6].
-Newman’s Lemma is used in its classic ARS form [10].
-For independence via termination, see Goodstein and Kirby-Paris and ordinal tools due to Buchholz [7, 8, 3].
-For fixed-target reachability complexity and modularity results, see [1, 4]; for undecidability at shallow depth in non-linear systems, see [9].
-Appendix B Conclusion
-This paper presents the KO7 operator-only kernel, proving strong normalization and confluence for the guarded safe subrelation.
-These results yield a certified normalizer and unique normal forms via Newman’s Lemma.
-An explicit non-local-join witness at the eqW reflexive peak (under $\kappa^{M}=0$) is established, blocking confluence for the full relation and motivating the SafeStep fragment.
-The decidability result explains the clash between single-level Gödel encodings and global termination.
-The impossibility of additive bumps under duplication is formally verified.
-Future directions include completing a stratified meta-layer, developing arithmetic encodings within KO7, and mapping the boundary between internally definable measures and external ordinal strength.

### v4

@@
-Strong Normalization for the Safe Fragment of a Minimal Rewrite System: A Triple-Lexicographic Proof and a Conjecture on the Unprovability of Full Termination for Any Relational Operator-Only TRS
+Strong Normalization for the Safe Fragment of a Minimal Rewrite Calculus: A Triple-Lexicographic Proof and a Conjecture on Full-Termination Limits for Pure Recursive Calculi (PRC)
 Abstract
-This paper presents a minimal operator-only term rewriting system (KO7) designed to investigate the foundations of self-verifying computation.
-The system consists of seven constructors and eight reduction rules, with no external axioms, no imported logic, and no pre-existing arithmetic.
-Standard termination methods fail on this system.
-We document ten distinct proof strategies that succumb to the system’s structural complexity: additive counters, simple lexicographic measures, polynomial interpretations, single-bit flags, pure ordinal measures, structural depth counters, naive multiset orderings, local helper lemmas, complex hybrid measures, and unchecked recursion.
-These methods collapse at the same barrier: the duplication of the step argument in the recursor.
-The rec-succ rule redistributes its argument into both the function application and the recursive call ($s\to f(s,\dots)$), an operation that algebraically defeats every additive measure ($M_{\text{after}}\geq M_{\text{before}}$ whenever $M_{s}\geq 1$).
-This failure is not a defect of the proof search but a necessary feature of any system capable of sequential computation.
+This paper introduces the concept of a Pure Recursive Calculus (PRC), a minimal computational substrate defined by four strict structural conditions: no variable binding, no external memory, no imported axioms, and an unrestricted step-duplicating recursor.
+In our PRC framing, this recursor is the mechanism used to represent sequential iteration.
+We present KO7, a concrete PRC consisting of seven constructors and eight reduction rules, designed to investigate the foundations of self-verifying computation.
+Standard termination methods fail on this calculus due to a core structural barrier: the duplication of the step argument in the recursor.
+The rec-succ rule redistributes its argument into both the function application and the recursive call ($s\to f(s,\dots)$), an operation that algebraically defeats additive measures ($M_{\text{after}}\geq M_{\text{before}}$ whenever $M_{s}\geq 1$).
+Because of this duplication, thirteen distinct internal proof strategies succumb to the calculus’s complexity: additive counters, simple lexicographic measures, polynomial interpretations, single-bit flags, pure ordinal measures, structural depth counters, naive multiset orderings, local helper lemmas, complex hybrid measures, unchecked recursion, pure precedence orders, linear weight functions (KBO-style), and standard tree depth.
+Furthermore, classical path orders (LPO/MPO) bypass this barrier by relying on the universal Subterm Property, which acts as an imported external axiom.
+Counting these external path-order methods yields fourteen total strategy families (Section 7).
+This pattern indicates a structural limitation for calculi of this form rather than a defect of proof search.
 To resolve this, we define SafeStep, a guarded fragment restricted by a $\delta$-phase bit.
-For this fragment, a novel triple-lexicographic measure (combining phase bits, Dershowitz-Manna multiset ordering, and ordinal ranking) succeeds, supporting a mechanically verified normalizer with proven totality and soundness.
-We state a fundamental conjecture: No relational operator-only TRS can have its full-system termination proved by internally definable methods.
-We define a “relational” TRS as one capable of internalized sequential computation.
-This property is shared by Turing machines (via transition functions), lambda calculus (via fixed-point combinators), and recursive functions.
-All such systems require a recursor to define the “next” step; this recursor inherently creates the duplication barrier.
-This result reveals a form of Pre-Arithmetic Incompleteness.
-Unlike Gödel’s theorems, which require arithmetic sufficient for encoding (Robinson Q), this barrier appears at a more primitive layer: in operator systems possessing only order and recursion, before arithmetic emerges.
-The system’s inability to prove its own termination from within is a structural limitation of ordered computation itself, paralleling the halting problem but manifesting at the substrate of operator definitions.
-The complete Lean 4 formalization ( 7,000 lines) is available at https://github.com/MosesRahnama/OperatorKO7.
+For this fragment, a triple-lexicographic measure (combining phase bits, Dershowitz-Manna multiset ordering, and a computable natural rank) succeeds, supporting a mechanically verified normalizer with proven totality and soundness.
+We state a central conjecture: For Pure Recursive Calculi (PRC), no internally definable method currently known to us proves full-calculus termination.
+Unlike Turing machines (which require an external tape), lambda calculus (which requires variable binding), and recursive functions (which import zero and successor as external primitives), a PRC has none of these external scaffolds.
+On this framing, the termination barrier is positioned at a more primitive layer than those classical settings.
+These findings motivate a form of Pre-Arithmetic Incompleteness.
+Unlike Gödel’s theorems, which require arithmetic sufficient for encoding (Robinson Q), the barrier studied here appears at a more primitive layer: in operator calculi possessing only order and recursion, before arithmetic emerges.
+In this setting, the difficulty of proving full termination from internal methods parallels halting-style limitations at the substrate of operator definitions.
+The complete Lean 4 formalization ($\sim$6,000 LOC) is available at https://github.com/MosesRahnama/OperatorKO7.
 1 Introduction
@@
 There are no binders, types, external axioms, or semantic predicates; the rules are the semantics.
-The primary goals are: (i) a clean, duplication-robust proof of strong normalization (SN); (ii) a certified normalizer that always returns a normal form; (iii) unique normal forms via Newman’s Lemma under a local-confluence assumption; and (iv) a conjecture regarding the unprovability of termination for relational operator-only systems using internal methods.
+The primary goals are: (i) a clean proof of strong normalization (SN) that handles duplication; (ii) a certified normalizer that always returns a normal form; (iii) unique normal forms via Newman’s Lemma, with the required local-confluence hypothesis discharged in the artifact; and (iv) a conjecture regarding the unprovability of termination for Pure Recursive Calculi (PRC) using internal methods.
 Scope.
 The formal results apply to a guarded safe subrelation.
-This restriction is necessary because the full relation is not locally confluent at the root peak $\mathrm{eqW}\,a\,a$ when $\kappa^{M}(a)=0$.
-The safe fragment employs guarded/context joiners to ensure unique normal forms (§6).
-A second conceptual goal is to situate KO7 against results about fixed-target reachability in terminating TRSs.
+This restriction is necessary because the full relation has an explicit non-local-join witness at $\mathrm{eqW}\;\mathrm{void}\;\mathrm{void}$.
+The safe fragment employs guarded/context joiners; the Newman local-join hypothesis is explicitly formulated and fully discharged in the artifact (§6).
+Why KO7?
+[…]
+A second conceptual goal is to situate KO7 against results about fixed-target reachability in terminating calculi.
 If an internal provability predicate is defined by “$t$ reduces to $\top$”, then under SN the set decidability of $\{t\mid t\to^{*}\top\}$ is reducible to normalization.
 With confluence, decision reduces to normal-form equality.
-This explains why single-level Gödel encodings cannot coexist with globally terminating proof search; stratification is required.
 Contributions
-Strong normalization is proven for a guarded fragment of KO7 using a novel triple-lexicographic measure that combines a phase bit, multiset ordering (Dershowitz-Manna), and ordinal ranking.
+Strong normalization is proven for a guarded fragment of KO7 using a triple-lexicographic measure that combines a phase bit, multiset ordering (Dershowitz-Manna), and a computable natural rank.
 This result supports a certified normalizer with proven totality and soundness.
-Furthermore, confluence for the safe fragment is established via Newman’s Lemma given verified local confluence.
-The work also includes formally verified impossibility results for additive and polynomial measures under duplication, demonstrating fundamental limitations in termination proving for self-referential systems.
+Furthermore, a Newman engine for the safe fragment is formalized, and the global local-join hypothesis is fully discharged, yielding unconditional confluence and unique normal forms for the safe fragment.
+Furthermore, we apply established proof theory to classify the boundaries of internal termination proving.
+Drawing on the work of Ferreira and Zantema [8], Buchholz [3], Gallier [9], and Rathjen–Weiermann [12], we observe that classical simplification orders (like LPO and MPO) fall outside our internal-definability boundary for pure calculi.
+These methods achieve termination by introducing the universal Subterm Property ($f(t)>t$) as a precondition, which in proof-theoretic analyses is tied to Kruskal-style well-foundedness arguments and transfinite induction strength (e.g., $\Pi^{1}_{1}$-comprehension).
+Because these orders import external structural axioms and well-ordering principles, they are outside the internal method class used in this paper.
+While not a proof of general non-termination, the work includes formally verified impossibility results for fourteen distinct strategy families under duplication.
+These machine-checked barriers provide evidence of structural limits in internal termination proving for self-referential calculi, directly motivating our central conjecture.
 Highlights (formalization summary)
-  • SN (SafeStep) via triple lex: Formally proven using a lexicographic measure combining a $\delta$-phase bit, a Dershowitz-Manna multiset rank, and an ordinal.
+  • SN (SafeStep) via triple lex: Formally proven using a lexicographic measure combining a $\delta$-phase bit, a Dershowitz-Manna multiset rank, and a computable natural rank.
   • Certified normalizer: A total and sound normalization function is defined by well-founded recursion.
-  • Newman (SafeStep): Confluence is established via Newman’s Lemma using a verified local confluence property for the safe fragment.
-  • Full Step caveat: A specific peak ($\mathrm{eqW}\,a\,a$ with $\kappa^{M}(a)=0$) where local join fails is exhibited, justifying the restriction to the safe subrelation.
-  • Impossibility results: The failure of simpler additive and polynomial measures is formally witnessed by counterexamples.
-2 Background: TRSs, SN, reachability, and Newman
+  • Newman (SafeStep): A Newman theorem is formalized for SafeStep; the global local-join hypothesis is discharged in the artifact, yielding unconditional confluence and unique normal forms for the safe fragment.
+  • Full Step caveat: A concrete non-local-join witness at $\mathrm{eqW}\;\mathrm{void}\;\mathrm{void}$ is exhibited for the full relation, justifying restriction to a guarded fragment.
+  • Impossibility results: Twelve distinct impossibility barriers (including additive, polynomial, weight-based, and precedence-based measures) are formally machine-checked in Impossibility_Lemmas.lean and Conjecture_Boundary.lean, alongside two meta-theoretical barriers.
+2 Background: Term Rewriting, SN, reachability, and Newman
 Standard abstract reduction and term rewriting notions apply [2, 13].
 A term is in normal form if no rule applies.
-A TRS is strongly normalizing (SN) if there are no infinite reductions.
+A rewrite calculus is strongly normalizing (SN) if there are no infinite reductions.
 A relation is confluent if for any $t\Rightarrow^{\ast}u$ and $t\Rightarrow^{\ast}v$ there exists $w$ with $u\Rightarrow^{\ast}w$ and $v\Rightarrow^{\ast}w$.
@@
 Newman’s Lemma asserts SN + local confluence $\Rightarrow$ confluence [11], yielding unique normal forms.
-Fixed-target (“small-term”) reachability in terminating TRSs has well-charted complexity: NP-complete for length-reducing systems (dropping to P under confluence), NExpTime/N2ExpTime for (linear) polynomial interpretations, and PSPACE for KBO-terminating TRSs [1].
-Modularity holds in certain linear, non-collapsing combinations [3], while even flat non-linear systems exhibit undecidable reachability and confluence [6].
+Fixed-target (“small-term”) reachability in terminating calculi has well-charted complexity: NP-complete for length-reducing calculi (dropping to P under confluence), NExpTime/N2ExpTime for (linear) polynomial interpretations, and PSPACE for KBO-terminating calculi [1].
+Modularity holds in certain linear, non-collapsing combinations [4], while even flat non-linear calculi exhibit undecidable reachability and confluence [10].
 Termination under duplication typically requires orders beyond plain sizes; a naive size can increase under duplicating rules.
@@
 3 The KO7 calculus
-KO7 is a finite TRS over a small signature (7 constructors) with 8 rules (including the equality-witness operator eqW). For concreteness we list the kernel rule shapes below.
-Table 1: KO7’s 8 kernel rules. In the full kernel relation, R_eq_refl and R_eq_diff overlap at eqW a a. The certified safe fragment guards R_eq_diff by $a\neq b$. R_merge_cancel is collapsing.
+KO7 is a finite rewrite calculus over a small signature (7 constructors) with 8 rules (including the equality-witness operator eqW). For concreteness we list the kernel rule shapes below.
+Table 1: KO7’s 8 kernel rules.
+In the full kernel relation, R_eq_refl and R_eq_diff overlap at eqW a a.
+The certified SafeStep fragment adds guards: R_eq_diff requires $a\neq b$, R_merge_void_left/right require $\delta$-flag tie conditions, R_merge_cancel requires both $\delta$-flag and $\kappa^{M}$ guards, R_rec_zero requires a $\delta$-flag guard on $b$, and R_eq_refl requires $\kappa^{M}(a)=0$.
+R_merge_cancel is collapsing.
+(* Note: R_rec_succ duplicates the step argument $s$ (one occurrence on the LHS, two on the RHS).
+This duplication is the termination stressor: the paper proves strong normalization for the guarded SafeStep fragment, not for the unrestricted full relation.)
 Rule Head Arity Shape Dup?
@@
   • Integrate/delta: $\mathrm{integrate}(\delta\,t)\Rightarrow^{\ast}\mathrm{void}$ (verified).
-  • Equality (meta-level consequence under confluence):
-    • If $\mathrm{nf}(a)=\mathrm{nf}(b)$ then $\mathrm{eqW}\,a\,b\Rightarrow^{\ast}\mathrm{void}$.
-    • Otherwise $\mathrm{eqW}\,a\,b\Rightarrow^{\ast}\mathrm{integrate}(\mathrm{merge}\,a\,b)$.
-    Note: In the safe fragment, confluence ensures these outcomes are unique.
+  • Strict Syntactic Equality:
+    • If $a$ and $b$ are syntactically identical ($a=b$), then $\mathrm{eqW}\,a\,b\Rightarrow\mathrm{void}$ (provided $\kappa^{M}(a)=0$).
+    • If $a\neq b$, then $\mathrm{eqW}\,a\,b\Rightarrow\mathrm{integrate}(\mathrm{merge}\,a\,b)$.
+    Note: In the safe fragment, arguments inside $\mathrm{eqW}$ are not subject to context closure, enforcing strict syntactic comparison rather than normal-form evaluation.
 Note. Strong normalization and Newman-based confluence are proven for the safe fragment.
 4 Strong normalization
-The termination argument relies on a triple-lexicographic measure
-$\mu^{3}(t):=(\delta\text{-flag}(t),\;\kappa^{M}(t),\;\mu_{\mathrm{ord}}(t))$
-ordered by the lex product of: (i) a phase bit dropping on the successor recursion; (ii) a multiset of ranks $\kappa^{M}$ (Dershowitz-Manna) with an explicit precedence/status orienting redex >> pieces; and (iii) an ordinal payload $\mu_{\mathrm{ord}}$ for non-duplicating ties.
-Ordinal hazards are formally handled (right-addition is not strictly monotone; absorption $\alpha+\beta=\beta$ requires $\omega\leq\beta$).
-Duplicating branches use a compact MPO-style head precedence.
-Step vs SafeStep (measure unification)
-For the SafeStep relation, a single unified triple lex order $(\delta,\kappa^{M},\mu_{\mathrm{ord}})$ suffices.
-Steps in the full relation are covered by a disjunctive decrease certificate: each step corresponds to either a KO7 lex drop or an MPO triple drop.
-Theorem 1 (Per-step decrease (SafeStep)).
-For every rule instance $t\Rightarrow t^{\prime}$ in the guarded SafeStep relation, we have $\mu^{3}(t^{\prime})<_{\mathrm{Lex}}\mu^{3}(t)$.
+The canonical termination argument relies on a computable triple-lexicographic measure
+$\mu^{3}_{c}(t):=(\delta\text{-flag}(t),\;\kappa^{M}(t),\;\tau(t))$
+ordered by the lex product of: (i) a phase bit dropping on the successor recursion; (ii) a multiset of ranks $\kappa^{M}$ (Dershowitz-Manna); and (iii) a computable natural payload $\tau$ for non-duplicating ties.
+Step vs SafeStep (artifact scope)
+For the SafeStep relation, a single unified computable triple lex order $(\delta,\kappa^{M},\tau)$ suffices and is the canonical certificate used in this paper (Meta/ComputableMeasure.lean, with core relation definitions in Meta/SafeStep_Core.lean).
+For the unrestricted full relation Step, termination is not claimed in this artifact.
+Theorem 4.1 (Per-step decrease (SafeStep)).
+For every rule instance $t\Rightarrow t^{\prime}$ in the guarded SafeStep relation, we have $\mu^{3}_{c}(t^{\prime})<_{\mathrm{Lex}}\mu^{3}_{c}(t)$.
 […]
-DM vs MPO on rules (explicit)
-For merge-cancel, we use a DM multiset lift over a base order where $\mathrm{merge}\,t\,t$ strictly dominates $t$.
-For eqW, we use a compact MPO-leaning measure where the head precedence orients $\mathrm{eqW}\,a\,b$ strictly above $\mathrm{integrate}(\mathrm{merge}\,a\,b)$.
-  • Base-order premise (merge-cancel): in $\mathrm{merge}\,t\,t\to t$, the RHS is strictly smaller than the LHS.
-  • Base-order premise (eqW): in $\mathrm{eqW}\,a\,b\to\mathrm{integrate}(\mathrm{merge}\,a\,b)$, the RHS is strictly smaller under head precedence.
+DM and $\tau$ roles on rules (explicit)
+For rec-zero, the $\kappa^{M}$ component provides a strict DM drop.
+For merge-cancel, the $\kappa^{M}$ component ties at 0, and the strict decrease is provided by the $\tau$ component.
+For rec-succ, the outer $\delta\text{-flag}$ drops from $1$ to 0.
+For eqW-rules in SafeStep (where $\kappa^{M}(a)=0$ is a guard), strict decrease is discharged in the computable $\tau$ component; the more general unguarded lemma in the artifact also includes a $\kappa^{M}(a)\neq 0$ branch discharged by DM.
+  • DM premise: in duplicating-sensitive branches, the multiset of RHS pieces is strictly DM-smaller than the removed LHS redex profile.
+  • $\tau$ premise: when the $\delta\text{-flag}$ and $\kappa^{M}$ tie, a direct natural-number inequality closes the branch.
 No-Go for constant bumps on $\kappa$ (generic duplicator)
-Lemma 1 (No fixed +k+k or boolean flag orients a generic duplicator).
+Lemma 4.2 (No fixed +k+k or boolean flag orients a generic duplicator).
 Let $\kappa$ be a max-depth-style counter and fix any $k\in\mathbb{N}$.
@@
 $\square$
-Remark 1.
+Remark 4.3.
 […]
@@
 so there is no strict drop when $\rho(S)\geq 1$.
-The robust fix uses DM/MPO: replace one element by a multiset of strictly smaller elements (DM), or use RPO/MPO with a precedence/status such that the LHS redex strictly dominates each RHS piece.
-Corollary 1 (Strong normalization).
+The fix used here is the computable triple stack $(\delta\text{-flag},\kappa^{M},\tau)$: DM handles multiset-sensitive branches and $\tau$ resolves ties.
+Corollary 4.4 (Strong normalization).
 The guarded relation SafeStep for KO7 is strongly normalizing.
@@
 Minimal counterpatterns that motivate the multiset/path components are recorded:
-  • Pure ordinal ($\mu$) only: shape-blind bounds fail to separate nested $\delta$ from its context.
-  • Additive bumps on $\kappa$: ties persist on duplicators (Lemma 1).
+  • Pure ordinal ($\mu$) only: shape-blind bounds fail to separate nested $\mathrm{delta}$ structures from their contexts.
+  • Additive bumps on $\kappa$: ties persist on duplicators (Lemma 4.2).
   • Additive counters $\rho$: by the identity $\rho(\text{after})=\rho(\text{before})-1+\rho(S)$ there is no strict drop when $\rho(S)\geq 1$.
   • Polynomial interpretations that rely on imported arithmetic or hard-coded constants: the duplicating rules force case splits that cannot be oriented without stepping outside the operator language, so these methods fail within the allowed internal toolbox.
-The fix is DM/MPO: ensure each RHS piece is strictly smaller than the removed redex in a base order and lift via a multiset/path extension.
+The final artifact path is the computable lex stack $\mu^{3}_{c}=(\delta,\kappa^{M},\tau)$, where DM handles multiset-sensitive branches and $\tau$ resolves tie cases.
 5 A certified normalizer
-A normalization function is defined by well-founded recursion on $\mu^{3}$. Properties proven in the formalization include:
-(Totality) $\forall t\,\exists n.~\mathrm{Normalize}(t)=n$
-(Soundness) $\forall t.~\mathrm{normal}(\mathrm{Normalize}(t))\;\wedge\;t\Rightarrow^{\ast}\,\mathrm{Normalize}(t).$
+A normalization function is defined by well-founded recursion on SafeStepRev, with well-foundedness certified by the computable measure stack $\mu^{3}_{c}$.
+The normalizer operates on the root step relation SafeStep (not the context closure SafeStepCtx); a term is in normal form when no root safe rule applies ($\mathrm{NormalFormSafe}(t)\iff\neg\exists u.\;\mathrm{SafeStep}\;t\;u$).
+Properties proven in the formalization include:
+(Totality) $\forall t\,\exists n.~\mathrm{SafeStepStar}(t,n)\land\mathrm{NormalFormSafe}(n)$
+(Soundness) $\forall t.~\mathrm{SafeStepStar}(t,\mathrm{normalizeSafe}(t))\land\mathrm{NormalFormSafe}(\mathrm{normalizeSafe}(t)).$
 No specific efficiency claim is made: worst-case normalization cost follows the termination witness (cf.
@@
 6 Local confluence and Newman (guarded safe relation)
-Local confluence is discharged by joining the finite set of critical pairs. Combining Cor. 1 with Newman’s Lemma [7] yields:
-Theorem 2 (Confluence and unique normal forms).
+The Newman engine is formalized and the global local-join hypothesis ($\mathsf{locAll}:\forall\,a,\,\mathrm{LocalJoinAt}\;a$) is fully discharged in the artifact.
+Combined with Cor.
+4.4, Newman’s Lemma [11] yields:
+Theorem 6.1 (Confluence and unique normal forms).
 If a relation is strongly normalizing and locally confluent, then it is confluent. Hence every term reduces to a unique normal form.
 Instantiation (SafeStep).
-Combining Cor.
-1 (SN for SafeStep) with local-join lemmas yields confluence and unique normal forms for the safe fragment by Newman’s Lemma.
-In contrast, the full relation is not locally confluent at the root peak $\mathrm{eqW}\,a\,a$ under $\kappa^{M}(a)=0$; thus full confluence does not hold.
-The SafeStep relation restricts eqW to cases where arguments are guarded, preventing this divergence.
+The formalization provides a comprehensive library of local-join lemmas over safe root/context shapes, fully discharging the $\mathsf{locAll}$ hypothesis to prove unconditional confluence.
+In contrast, the full relation has an explicit non-local-join witness at $\mathrm{eqW}\;\mathrm{void}\;\mathrm{void}$; thus unrestricted full-step confluence is not claimed.
+The SafeStep relation introduces guard conditions that avoid this divergence pattern.
 The star-star join proof follows the standard accessibility (Acc) recursion at the source, with a case split over star shapes (“head step + tail”) and composition via transitivity.
@@
 The formal results concern a guarded safe subrelation SafeStep for which we prove:
-  • SN (SafeStep). The KO7 triple measure $\mu^{3}$ strictly drops on every SafeStep. This yields a certified normalizer that is total and sound for the safe fragment.
-  • Local Confluence (SafeStep). Local-join lemmas per root shape and context wrappers are provided. Newman then yields confluence and unique NFs for the safe fragment.
+  • SN (SafeStep). The computable KO7 triple measure $\mu^{3}_{c}$ strictly drops on every SafeStep. This yields a certified normalizer that is total and sound for the safe fragment.
+  • Local-Join Library (SafeStep).
+Local-join lemmas per root shape and context wrappers are provided; the global hypothesis $\mathsf{locAll}$ is discharged in the artifact (locAll_safe), yielding unconditional confluence and unique NFs.
 Critical-pair coverage
-The following table covers the safe root configurations with explicit local-join lemmas.
-This is exhaustive for SafeStep except the reflexive $\mathrm{eqW}\,a\,a$ peak under $\kappa^{M}(a)=0$, which is shown to be not locally joinable at the root.
-We discharge many eqW cases via guarded/context wrappers.
+The following table records representative safe root configurations with explicit local-join lemmas.
+We also exhibit a full-step non-local-join witness at $\mathrm{eqW}\;\mathrm{void}\;\mathrm{void}$ and discharge many guarded eqW cases via context wrappers.
+Guarded variants exclude spurious branches.
+Contextual wrappers lift root joins to context.
 Table 2: Critical-pair coverage at the root for SafeStep.
@@
 $\mathrm{eqW}\,a\,b,~a\neq b$ localJoin_eqW_ne (unique target $\mathrm{integrate}(\mathrm{merge}\,a\,b)$)
-$\mathrm{eqW}\,a\,a,~\kappa^{M}(a)=0$ not locally joinable at root
-Guarded variants exclude spurious branches. Contextual wrappers lift root joins to context.
+$\mathrm{eqW}\,a\,a,~\kappa^{M}(a)=0$ localJoin_all_safe (eqW/refl branch; target void)
+For the $\mathrm{eqW}\,a\,a$ row, the relevant branch inside localJoin_all_safe is discharged via localJoin_of_unique.
 $\delta$-guard: definition and decidability
 Define the safe-phase predicate by $\delta\text{-guard}(t)\iff\mathrm{deltaFlag}(t)=0$.
-Here $\mathrm{deltaFlag}:\mathrm{Term}\to\mathbb{N}$ is a structurally recursive function defined on terms in the artifact (tracking split parity), so the predicate is decidable.
+Here $\mathrm{deltaFlag}:\mathrm{Trace}\to\mathbb{N}$ is a binary root-shape detector for $\mathrm{rec}\Delta\,\_\,\_\,(\mathrm{delta}\,\_)$ defined in the artifact, so the predicate is decidable.
 Facts: $\mathrm{deltaFlag}(\mathrm{eqW}\,a\,b)=0$; merge-void rules require $\mathrm{deltaFlag}(t)=0$.
 […]
-$\mathrm{rec}\_\Delta\,b\,s\,(\delta\,n)$$\mathrm{app}\;s\;(\mathrm{rec}\_\Delta\,b\,s\,n)$$\delta$-flag drop
-7 Impossibility results: The Ten Failed Methods
-The formal development establishes the failure of ten distinct termination strategies, necessitating the use of the triple-lexicographic measure with Dershowitz-Manna orderings.
-These failures are not merely heuristic; they are formally witnessed by counterexamples in the codebase (specifically in Meta/Impossibility_Lemmas.lean and Meta/Operational_Incompleteness.lean).
+$\mathrm{rec}\Delta\,b\,s\,(\mathrm{delta}\,n)$$\mathrm{app}\;s\;(\mathrm{rec}\Delta\,b\,s\,n)$$\delta\text{-flag}$ drop
+7 Impossibility results: The Fourteen Failed Methods
+We document the failure of fourteen distinct termination strategy families, motivating the triple-lexicographic measure with Dershowitz-Manna orderings used in the safe fragment.
+Twelve of these failures are not merely heuristic; they are formally witnessed by machine-checked counterexamples in the codebase (specifically in Meta/Impossibility_Lemmas.lean and Meta/Conjecture_Boundary.lean).
+The remaining two (local helper lemmas and path orders) are treated as meta-theoretical and methodological barriers.
   1. 1.
@@
   4. 4.
-    Single-Bit Flags: A boolean flag alone is non-monotone; specific steps (e.g., ‘merge void‘) can toggle the flag from 0 to 1.
+    Single-Bit Flags: A boolean flag alone is non-monotone; specific steps (e.g., ‘merge void’) can toggle the flag from 0 to 1.
   5. 5.
@@
   6. 6.
-    Structural Depth: Simple depth counters ($\kappa$) fail on collapsing rules like merge_cancel where the result may have equal depth to the input.
+    Structural Depth: Simple depth counters ($\kappa$) fail on collapsing rules like ‘merge_cancel’ where the result may have equal depth to the input.
   7. 7.
@@
   8. 8.
-    Local Helper Lemmas: Attempts to patch specific rules with local lemmas failed due to the global interconnectedness of the recursion.
+    Local Helper Lemmas (Methodological): Attempts to patch specific rules with local lemmas failed due to the global interconnectedness of the recursion.
   9. 9.
@@
   10. 10.
-    Unchecked Recursion: The raw rec_succ rule itself serves as the ultimate counterexample; without the SafeStep guards, it permits unbounded recursion that no internal measure can bound.
-These ten failures collectively demonstrate that the termination barrier is structural: the system’s internal logic cannot capture the necessary invariants without the sophisticated machinery of the triple-lexicographic measure.
+    Unchecked Recursion: The raw ‘rec_succ’ rule itself serves as a counterexample pattern; without the SafeStep guards, it permits unbounded recursion that is not handled by the audited internal families.
+  11. 11.
+    Path Orders (Meta-Theoretical): Rely on the universal Subterm Property ($f(t)>t$), which injects an external structural axiom and artificial fixed point not present in the pure calculus.
+  12. 12.
+    Pure Precedence Orders: Measures relying strictly on head constructor precedence fail on collapsing rules like merge_cancel because the RHS can have the exact same head constructor as the LHS.
+  13. 13.
+    Linear Weight Functions (KBO-style): Purely additive weight functions fail on the duplication in rec_succ regardless of the constants chosen, as the duplicated argument can grow arbitrarily large.
+  14. 14.
+    Standard Tree Depth: Measuring the maximum depth of the Abstract Syntax Tree (AST) fails because the duplication of $s$ in rec_succ can strictly increase the overall depth of the term.
+Taken together, these fourteen failures provide evidence that the termination barrier is structural: within the audited internal families, the necessary invariants are not captured without the triple-lexicographic measure used for SafeStep.
+The Boundary Between Code and Meta-Theory
+We distinguish between what is machine-checked in the code and what is argued in the meta-theory regarding Path Orders (Item 11).
+The Lean artifact (Conjecture_Boundary.lean) formally demonstrates that the isolated components of path orders (pure precedence, Item 12, and linear weights, Item 13) mathematically fail to orient the calculus.
+In classical proofs, successful path-order arguments rely on the universal Subterm Property ($f(t)>t$).
+While full LPO is a classical external method that could theoretically orient such systems, doing so imports this property.
+As argued in Section 8.2, this is a meta-theoretical intervention (via Kruskal-style assumptions) rather than an internally definable proof under Definition 9.1.
+The code demonstrates pressure toward such an axiom; the text analyzes its scope.
 Polynomial Impossibility
@@
 This fails for unbounded $M(x)$.
-More generally: any polynomial proof for a system with duplicating rules must hardcode specific constants (e.g., $M(\mathrm{void})=2$) to satisfy inequalities like $2M(s)>M(s)+1$.
+More generally for the duplication pattern studied here: polynomial proofs typically require hardcoded constants (e.g., $M(\mathrm{void})=2$) to satisfy inequalities like $2M(s)>M(s)+1$.
 Such proofs import external arithmetic properties and impose arbitrary values on operators, violating the principle that operators’ semantics should be defined solely by rewrite rules.
-Changing the hardcoded constant (e.g., $M(\mathrm{void})=1$) collapses the proof.
-This structural failure is why polynomial interpretations cannot orient generic duplicating rules without external axioms.
-Internally definable measures
-A simple contract for internally defined termination measures structures these negative results:
-Definition 1 (Internally definable measure).
-An internally definable measure for a type $\alpha$ consists of $(\beta,\;<\_{\beta},\;\mathsf{wf},\;m,\;\mathsf{ctxMono},\;\mathsf{piecesLt})$ where: $\beta$ is a base order carrier; $_{\beta}$ is a well-founded relation with explicit witness $\mathsf{wf}$; $m:\alpha\to\beta$ is the measure; $\mathsf{ctxMono}$ expresses context compatibility; and $\mathsf{piecesLt}$ asserts that in each rule instance, every RHS piece is strictly smaller than the removed LHS redex w.r.t.
-$_{\beta}$.
+Changing the hardcoded constant (e.g., $M(\mathrm{void})=1$) collapses the orientation argument.
+This is why polynomial interpretations in this setting are treated as outside the intended internal method class.
 8 Decidability of Reachability
-Theorem 3 (Fixed-target reachability).
-In a strongly normalizing TRS, the set $\{t\mid t\Rightarrow^{*}c\}$ for any constant $c$ is decidable via normalization.
+Theorem 8.1 (Fixed-target reachability).
+In a strongly normalizing and confluent calculus, the set $\{t\mid t\Rightarrow^{*}c\}$ for any constant $c$ is decidable via normalization.
 This connects the work to classical results: encoding undecidable properties through reduction to constants violates known theoretical limits.
@@
 Assumptions (model)
-The model assumes a finite first-order TRS over finite terms.
-The decision procedure computes a normal form (by SN) and checks whether it is $\top$; with local confluence, this reduces to one normalization and an equality test.
+The model assumes a finite first-order rewrite calculus over finite terms.
+The decision procedure computes a normal form (by SN) and checks whether it is $\top$; with confluence (hence unique normal forms), this reduces to one normalization and an equality test.
 Complexity context
-Small-term reachability in terminating TRSs ranges from NP (length-reducing) to NExpTime/N2ExpTime (polynomial interpretations) and PSPACE (KBO), with confluence lowering the length-reducing class to P [1].
-This situates our “normalize and compare” decision procedure for KO7 within the established landscape.
-Moreover, decidability can be modular for disjoint unions under left-linearity/non-collapsing assumptions [3], while termination alone does not guarantee decidability: even flat non-linear TRSs have undecidable reachability/joinability/confluence [6].
-9 Conjecture: Full Termination of Any Relational Operator-Only TRS
-What “relational” means
-A relational operator-only TRS is one capable of sequential computation: any system that can express “do $X$, then do YY” without external conditions.
-The terms are equivalent: “relational” = “capable of sequential computation.”
-9.1 The Relational Nature of KO7 and Recursors
-The classification of KO7 as “relational” is not arbitrary; it describes the system’s capability to internalize sequential control flow.
-In flat rewrite systems (e.g., group axioms), rules rewrite terms to equivalent terms.
-In a relational TRS like KO7, the rewrite rules define a computational process.
+Small-term reachability in terminating calculi ranges from NP (length-reducing) to NExpTime/N2ExpTime (polynomial interpretations) and PSPACE (KBO), with confluence lowering the length-reducing class to P [1].
+This situates our “normalize and compare” decision procedure for KO7 within the known complexity results.
+Moreover, decidability can be modular for disjoint unions under left-linearity/non-collapsing assumptions [4], while termination alone does not guarantee decidability: even flat non-linear calculi have undecidable reachability/joinability/confluence [10].
+9 Conjecture: Full Termination Limits in Pure Recursive Calculi (PRC)
+Sequential Computation
+A Pure Recursive Calculus is one capable of sequential computation: any calculus that can express “do $X$, then do YY” without external conditions.
+9.1 The Sequential Nature of KO7 and Recursors
+The classification of KO7 as a PRC is not arbitrary; it describes the calculus’s capability to internalize sequential control flow.
+In flat rewrite calculi (e.g., group axioms), rules rewrite terms to equivalent terms.
+In a sequential calculus like KO7, the rewrite rules define a computational process.
 […]
@@
 It applies the “step” term $s$ repeatedly, $n$ times.
-This is functionally identical to the transition function $\delta$ of a Turing machine or the fixed-point combinator $Y$ in lambda calculus.
-Both mechanisms allow a system to execute a sequence of operations of unbounded length.
+This is functionally identical to the recursor in Gödel’s System T or the step function of primitive recursive arithmetic.
+Both mechanisms allow a calculus to execute a sequence of operations of unbounded length.
 The Recursor Trap
-Any system capable of this sequential computation must have a mechanism to carry the “step” logic ($s$) forward into the next iteration.
+Within this PRC framing, sequential computation requires a mechanism to carry the “step” logic ($s$) into the next iteration.
 In KO7, this is achieved by duplicating $s$: it appears once in the current application ($\mathrm{app}\,s\dots$) and once in the recursive call ($\mathrm{rec}\Delta\dots s\dots$).
 This duplication is the structural barrier.
-As demonstrated in the Impossibility Results (Section 7), this redistribution of $s$ defeats all additive termination measures ($M_{\text{after}}\geq M_{\text{before}}$ whenever $M_{s}\geq 1$).
-A system that internalizes its own control flow in this way cannot prove its own termination using simple internal metrics, because the metric must account for the unbounded potential of the duplicated step function.
+As demonstrated in the Impossibility Results (Section 7) and formally witnessed in the artifact by the theorem rec_succ_additive_barrier, this redistribution of $s$ defeats additive termination measures ($M_{\text{after}}\geq M_{\text{before}}$ whenever $M_{s}\geq 1$).
+A calculus that internalizes control flow in this way faces a sharp difficulty when proving full termination with simple internal metrics, because the metric must account for the unbounded potential of the duplicated step function.
 Escaping the Trap: Meta-Layers vs. Internal Proofs
-Standard systems escape this trap by stepping outside the system.
+Standard calculi escape this trap by stepping outside the calculus.
   • Proof Assistants (Coq, Lean): Use a meta-logical layer (guard predicates, structural recursion checks) to enforce termination externally.
-The system doesn’t prove itself; the compiler proves the system.
+The object calculus does not prove termination by itself; the meta-system checks the required obligations.
   • Turing Machines: We do not ask the tape to prove it halts; we analyze the machine from the outside (and finding that we cannot, we discover the Halting Problem).
-KO7’s SafeStep fragment follows this pattern by imposing external guards ($\delta$-flag). The Conjecture posits that for the full, unguarded system, no such internal proof exists.
+KO7’s SafeStep fragment follows this pattern by imposing external guards ($\delta$-flag and $\kappa^{M}$ side conditions).
+The conjecture below is scoped to Pure Recursive Calculi (Definition 9.2), not to all operator-only kernels.
 The structural barrier
-Any operator system capable of ordered computation requires a recursor.
-A recursor redistributes its step argument across recursive calls.
+In this PRC setting, ordered computation is represented with a recursor.
+The recursor redistributes its step argument across recursive calls.
 This redistribution defeats additive termination measures: $M(\text{after})=M(\text{before})-1+M(s)$, yielding no strict drop when $M(s)\geq 1$.
-Definition 2 (Internally definable measure).
-An internally definable termination method for signature $\Sigma$ uses only: simplification orders with fixed precedence/status on $\Sigma$ (LPO/RPO/MPO), DM-multiset lifts of $\mathbb{N}$-valued ranks, and algebraic interpretations definable without importing external axioms.
+Definition 9.1 (Internally definable measure).
+An internally definable termination method for signature $\Sigma$ uses only: DM-multiset lifts of $\mathbb{N}$-valued ranks, and algebraic interpretations definable without importing external axioms.
 No encodings into external arithmetic, no borrowed logic, and no hard-coded constants outside the operator language.
-Conjecture 1 (Full Termination Conjecture for Relational TRSs).
-No relational operator-only TRS can have its full-system termination proved by internally definable methods. Specifically: let $R$ be an operator-only TRS with a recursor rule of the form
+Importantly, an internally definable method must be derived strictly from the operational semantics (the rewrite rules).
+It cannot inject universal structural axioms from the meta-theory.
+For example, standard simplification orders (LPO/RPO/MPO) are explicitly excluded because they rely on the universal Subterm Property ($f(t)>t$).
+As demonstrated by Ferreira and Zantema [8], this property acts as an imported precondition axiom, artificially establishing a fixed point (an anchor) that the pure calculus does not possess.
+Furthermore, as established in the proof-theoretic literature (e.g., Buchholz [3], Dershowitz & Okada [6]), the well-foundedness of these orders relies on Kruskal’s Tree Theorem, which Gallier [9] notes is a classic independence phenomenon requiring transfinite induction at or above $\varepsilon_{0}$.
+The distinction is: an internally definable method must be self-contained within the operator language; any proof that borrows structural axioms or well-ordering strength from an external theory is performing meta-level reasoning, not internal reasoning.
+Definition 9.2 (Pure Recursive Calculus (PRC)).
+A finite first-order rewrite calculus is a Pure Recursive Calculus (PRC) if it satisfies all four of the following conditions:
+  1. 1.
+    No variable binding. The signature contains no binders; there is no substitution meta-theory and no free variables. All term structure is encoded via constructors, not via variable names.
+  2. 2.
+    No external memory. The calculus has no external tape, registers, or environment. All computational state is carried within the term itself.
+  3. 3.
+    No imported axioms.
+All primitive objects are defined purely as syntactic constructors (e.g., void) governed by the reduction rules, rather than being imported as external mathematical constants with pre-existing axiomatic properties.
+  4. 4.
+    Unrestricted step-duplicating recursor. The signature contains a recursor rule of the form
+    $\mathrm{rec}(b,\,s,\,\sigma(n))\;\to\;f(s,\,\mathrm{rec}(b,\,s,\,n))$
+    in which the step argument $s$ appears on both sides and is unrestricted: no guard condition, type annotation, or structural bound is imposed on $s$.
+KO7 with full Step is a PRC.
+Classical Turing machines fail condition 2 (they require an external tape).
+The untyped lambda calculus fails condition 1 ($\lambda$-abstraction introduces variable binding).
+Systems of primitive recursive functions fail condition 3 (zero and successor are imported as external primitives, not operator-defined).
+In the sense of this definition, the class targets a more primitive setting than these classical computational models: the termination barrier identified here arises before arithmetic, variable binding, or external memory enter the picture.
+9.2 The Subterm Axiom and Artificial Fixed Points
+When classical rewriting theory encounters the duplication barrier, it typically deploys simplification orders (such as LPO or MPO).
+These orders succeed by enforcing the Subterm Property: the universal rule that any term is strictly greater than its subterms ($f(t)>t$).
+However, in the context of a Pure Recursive Calculus, this is a meta-theoretical intervention.
+While the subterm property is a natural consequence of the inductive definition of terms in the meta-theory, importing that inductive structure into the termination measure of a pure calculus constitutes an external intervention.
+The pure calculus only “knows” what its reduction rules explicitly define.
+It does not inherently know that $\mathrm{delta}(t)$ is greater than $t$ unless a rule reduces $\mathrm{delta}(t)$ to $t$.
+By forcing $f(t)>t$ universally across all constructors, simplification orders impose an external structural constraint on the system.
+They establish an artificial fixed point (an external anchor or “zero”) that the relational logic never defined.
+This suggests a structural limitation: proving termination of a pure sequential calculus may require introducing structural axioms at the meta-level.
+In this analysis, the anchor is introduced from outside the object calculus.
+This structural observation aligns with established proof theory.
+As demonstrated by Ferreira and Zantema [8], the well-foundedness of path orderings depends on assuming the subterm property as a precondition.
+Furthermore, in the proof-theoretic literature (e.g., Buchholz [3], Dershowitz & Okada [6]), these orders are formalized as inductively generated relations (least fixed points) whose well-foundedness relies on Kruskal’s Tree Theorem.
+From the perspective of reverse mathematics, Gallier [9] notes that Kruskal’s theorem is a classic independence phenomenon.
+As established by Rathjen and Weiermann [12], it requires proof-theoretic strength equivalent to $\Pi^{1}_{1}$-comprehension ($\Pi^{1}_{1}$-CA0) and ordinal bounds reaching into the Veblen hierarchy.
+Thus, in this framework, orienting a PRC with LPO is treated as meta-level reasoning rather than an internally definable termination proof.
+9.3 The Pre-Arithmetic Incompleteness Hierarchy
+The conjecture situates KO7 within a three-layer hierarchy of incompleteness results, presented as requiring progressively more external structure:
+Table 3: Three layers of computational incompleteness.
+PRC (Layer 0) requires only an object-level recursor and no external scaffolding; Turing machines (Layer 1) add an external tape and meta-level iteration; Gödel systems (Layer 2) additionally require arithmetic encoding.
+Layer Calculus class Minimum requirement Barrier
+0 (this work) PRC Order + object-level recursor No known internally definable proof of full termination
+1 (Turing) General computation External tape + meta-level transition Cannot decide halting (undecidable)
+2 (Gödel) Arithmetic calculus Robinson Q (++ and $\times$) Cannot prove own consistency internally
+Why Layer 0 Is Placed Below Layer 1
+A classical Turing machine performs sequential computation via a meta-level transition function $\delta:Q\times\Gamma\to Q\times\Gamma\times\{L,R\}$.
+The iteration (“run for $k$ steps”) is defined outside the machine as a sequence of configuration yields.
+The tape is external memory; there is no operator within the term language that internalizes the step.
+In a PRC, by contrast, the recursor $\mathrm{rec}\Delta$ is an object-level constructor: “do the next step” is a rewrite rule applied from inside the term, not a meta-operation on a configuration.
+This is why a Turing machine fails condition 2 of Definition 9.2 and falls outside the PRC class.
+Why Layer 0 Is Placed Below Layer 2
+Gödel’s incompleteness theorems require arithmetic sufficient for encoding, at minimum Robinson arithmetic $Q$, with addition, multiplication, zero, and successor as imported primitives.
+Presburger arithmetic (addition only) and Skolem arithmetic (multiplication only) are both decidable and complete; they escape Gödel because they lack the encoding power of full $Q$.
+A PRC requires none of this: no imported arithmetic, no Gödel numbering, no self-referential encoding.
+The termination barrier at Layer 0 arises from structural duplication in the recursor (a constructive phenomenon) rather than from a diagonal self-reference argument.
+Arithmetic itself is a higher-layer phenomenon: it emerges when the PRC operators are later interpreted (e.g., $\delta$ as successor, $\mathrm{rec}$ as primitive recursion).
+The Layer 0 barrier exists before that interpretation is imposed.
+This explains why single-level Gödel encodings cannot coexist with globally terminating proof search; stratification is required.
+Conjecture 9.3 (Scoped Full-Termination Conjecture for Pure Recursive Calculi).
+For Pure Recursive Calculi (Definition 9.2), no internally definable measure currently known to us proves full-calculus termination.
+Specifically: let $R$ be an operator-only calculus with a recursor rule of the form
 $\mathrm{rec}(b,s,\sigma(n))\to f(s,\mathrm{rec}(b,s,n)).$
-No internally definable measure (Definition 2) proves termination of $R$ when the step argument $s$ is unrestricted.
-The structural redistribution of $s$ into both the function application and the recursive call defeats every additive, polynomial, or path-order measure that does not import external arithmetic or ordinal axioms.
-Any proof that relies on external axioms, imported arithmetic, or meta-level encodings is outside scope and does not count as an internal termination argument.
+No internally definable measure (Definition 9.1) is currently known to prove termination of such an $R$ when the step argument $s$ is unrestricted. Specifically:
+  • Additive and polynomial methods are formally excluded in the audited family.
+The redistribution identity $M(\text{after})=M(\text{before})-1+M(s)$ yields no strict drop when $M(s)\geq 1$; polynomial interpretations without imported arithmetic degrade to additive measures and fail by the same identity (Lemma 4.2).
+Machine-checked barriers in Meta/Conjecture_Boundary.lean formally witness failure of additive counters, simple lexicographic pairs, flag-only methods, polynomial multiplicative ties, structural depth, and constellation-style hybrids (Section 7).
+  • The Subterm Axiom Boundary (LPO/RPO/MPO): Standard path orders are explicitly excluded from internally definable methods because they rely on the universal Subterm Property ($f(t)>t$).
+This property acts as an imported structural axiom, establishing an anchor that the pure calculus does not define via its own rewrite rules.
+Forcing $f(t)>t$ universally bypasses the duplication obstruction by adding meta-level structure.
+Under Definition 9.1, importing this axiom violates the “no imported axioms” boundary for PRCs; therefore, such arguments are treated as meta-theoretical rather than internal proofs.
+    Note on orientation layers: The canonical SafeStep termination proof in this paper is the computable stack $(\delta,\kappa^{M},\tau)$ from Meta/ComputableMeasure.lean: rec_zero is discharged by DM on $\kappa^{M}$, rec_succ by the $\delta$ drop, and merge_cancel/eqW tie branches by $\tau$.
+MPO/LPO appear only as external meta-theoretic comparators in the boundary discussion, not as the internal certified certificate.
+Any proof that relies on external axioms, imported arithmetic, or meta-level encodings is outside this scope.
 Scope of the conjecture
-The conjecture applies to any operator-only TRS capable of sequential computation: any system that can express “do $X$, then do YY” using only rewrite rules, without external axioms, imported arithmetic, or borrowed logic.
-Systems without such capability (purely ground, non-recursive, or lacking step-iteration) fall outside the conjecture’s scope.
-The claim is that once a system can compute sequentially without external conditions, its full termination escapes internal proof methods.
+The conjecture applies to Pure Recursive Calculi (Definition 9.2) under the internal-method boundary above.
+Calculi that fail any of the four PRC conditions (by introducing external memory, variable binding, imported axioms, or a guarded recursor) may lie outside this scope.
 Evidence
-In KO7, the rec-succ rule has the critical shape: the step argument $s$ appears twice on the RHS, and when $s$ contains nested redexes, every additive measure fails to decrease strictly.
+In KO7, the rec-succ rule has the critical shape: the step argument $s$ appears twice on the RHS, and when $s$ contains nested redexes, additive measures fail to decrease strictly.
 The redistribution identity $M(\text{after})=M(\text{before})-1+M(s)$ yields no strict drop when $M(s)\geq 1$.
-The guarded SafeStep fragment terminates because the $\delta$-phase bit restricts allowable step arguments; without such restriction, the full system admits unbounded recursive depth.
-The gap between internally definable ranking functions and what would be required cannot be closed without external methods (axioms, encodings, or borrowed arithmetic).
+The guarded SafeStep fragment terminates because the $\delta\text{-flag}$ restricts allowable step arguments; without such restriction, the full calculus admits unbounded recursive depth.
+At present, no internally definable ranking-function closure is known in this setting without external methods (axioms, encodings, or borrowed arithmetic).
 10 Formalization structure
@@
   • Impossibility results: Contains the verified counterexamples for additive measures and the proofs that simple measures fail under duplication.
-All claimed results are formally proven without reliance on unproven postulates in the current build.
-The formal development (Lean 4,  7,000 LOC; Apache-2.0 licensed) is available at https://github.com/MosesRahnama/OperatorKO7 and includes build instructions and the toolchain version.
+All theorem statements currently asserted in the build are formally proven without unproven postulates.
+Confluence theorems in Meta/Newman_Safe.lean are stated with a global local-join parameter ($\mathsf{locAll}$).
+The artifact discharges this parameter via locAll_safe from Meta/Confluence_Safe.lean to obtain unconditional confluentSafe and unique normal forms for SafeStep.
+The formal development (Lean 4, $\sim$6,000 LOC; dual-licensed for academic and commercial use) is available at https://github.com/MosesRahnama/OperatorKO7 and includes build instructions and the toolchain version.
 11 Conclusion
-This paper presents the KO7 operator-only kernel, proving strong normalization and confluence for the guarded safe subrelation.
-These results yield a certified normalizer and unique normal forms via Newman’s Lemma.
-An explicit non-local-join witness at the eqW reflexive peak (under $\kappa^{M}=0$) is established, blocking confluence for the full relation and motivating the SafeStep fragment.
+This paper presents the KO7 operator-only kernel, proving strong normalization for a guarded safe subrelation and formalizing a Newman confluence engine.
+These results yield a certified normalizer; unique normal forms follow unconditionally for the safe fragment.
+An explicit non-local-join witness at $\mathrm{eqW}\;\mathrm{void}\;\mathrm{void}$ is established for the full relation, motivating the SafeStep fragment.
 The decidability result explains the clash between single-level Gödel encodings and global termination.
 The impossibility of additive bumps under duplication is formally verified.
-Future directions include completing a stratified meta-layer, developing arithmetic encodings within KO7, and mapping the boundary between internally definable measures and external ordinal strength.
 Appendix A Addendum
@@
 The authoritative mechanized kernel is the Lean file OperatorKO7/Kernel.lean in the repository https://github.com/MosesRahnama/OperatorKO7. The safe fragment and proofs are under OperatorKO7/Meta/.
-A.2 Triple-lexicographic measure and duplication handling (DM/MPO)
-The measure $\mu^{3}(t)=(\delta\text{-flag}(t),\,\kappa^{M}(t),\,\mu_{\mathrm{ord}}(t))$ is ordered by lex:
+A.2 Auxiliary orientation sketch (DM/MPO)
+This subsection records historical orientation notes used for boundary analysis.
+It is not the canonical SafeStep termination certificate in this paper, which is the computable stack in Meta/ComputableMeasure.lean.
+An auxiliary triple $\widetilde{\mu}^{3}(t)=(\delta\text{-flag}(t),\,\kappa^{M}(t),\,r_{\mathrm{aux}}(t))$ is ordered lexicographically:
   • $\delta$-flag: phase bit dropping on the successor recursion branch.
   • Multiset $\kappa^{M}$: a Dershowitz-Manna multiset extension over a base precedence/status (or MPO head precedence) orienting redexes strictly above pieces; covers duplicators.
-  • Ordinal $\mu_{\mathrm{ord}}$: resolves non-duplicating ties; right-addition is not strictly monotone; absorption $\alpha+\beta=\beta$ needs $\omega\leq\beta$.
+  • Tie-break rank $r_{\mathrm{aux}}$: resolves non-duplicating ties when the first two coordinates are equal.
 Per-rule lemmas show every RHS component is strictly smaller than the removed redex in the base order.
-Aggregation for $\kappa^{M}$ (union, not sum)
-It is emphasized that $\kappa^{M}$ aggregates via multiset union ($\cup$), not numeric addition.
+Aggregation for $\kappa^{M}$
+It is emphasized that $\kappa^{M}$ aggregates via multiset operations (union $\cup$ and addition ++), not scalar numeric addition.
 For duplicating rules, the multiset of piece-weights on the RHS is DM-smaller than the singleton multiset containing the LHS redex weight, yielding a strict drop in the $\kappa^{M}$ component.
-For non-duplicating rules, $\kappa^{M}$ ties by definitional equality and the ordinal $\mu_{\mathrm{ord}}$ resolves the branch.
-In particular, for unguarded instances of eqW-refl and merge-cancel a $\kappa$-branch strategy is employed: when $\kappa^{M}(a)\neq 0$ a left-lex drop is obtained via DM; when $\kappa^{M}(a)=0$ the $\kappa^{M}$ component ties by rfl and the strict decrease is witnessed in the $\mu_{\mathrm{ord}}$ coordinate (right branch of Prod.Lex).
+For non-duplicating rules, $\kappa^{M}$ ties by definitional equality and $r_{\mathrm{aux}}$ resolves the branch.
+In particular, for unguarded instances of eqW-refl and merge-cancel a $\kappa$-branch strategy is employed: when $\kappa^{M}(a)\neq 0$ a left-lex drop is obtained via DM; when $\kappa^{M}(a)=0$ the $\kappa^{M}$ component ties by rfl and the strict decrease is witnessed in the $r_{\mathrm{aux}}$ coordinate (right branch of Prod.Lex).
 Short witness snippets (toy duplication)
@@
 […]
-Figure 1: Triple-lex measure components. Duplicators decrease via $\kappa^{M}$; non-duplicating ties via $\mu_{\mathrm{ord}}$.
+Figure 1: Auxiliary triple-lex components. Duplicators decrease via $\kappa^{M}$; non-duplicating ties via $r_{\mathrm{aux}}$.
 A.3 Newman scope (guarded safe relation)
@@
 Scope note: all confluence statements and their Newman instantiations are for SafeStep only.
-Full Step is not locally joinable at root for eqW a a with $\kappa^{M}(a)=0$; accordingly full-step confluence is not claimed.
+Full Step has an explicit non-local-join witness at root, eqW void void; accordingly full-step confluence is not claimed.
 A.4 Module map
@@
 Note: context closure and local-join lemmas are explicitly instantiated for the SafeStep fragment (SafeStepCtx) rather than the full relation.
-[…]
-Table 3: Per-rule decrease witnesses used in the SafeStep termination proof.
-Rule Base component Precedence/Status Witness Source
-merge void left $\mu$ (ordinal) (n/a) Theorem 4.1 Termination
-merge void right $\mu$ (ordinal) (n/a) Theorem 4.2 Termination
-merge cancel DM on $\kappa^{M}$ (redex >> pieces) head precedence Theorem 4.3 Termination
-rec zero DM on $\kappa^{M}$ rec >> pieces Theorem 4.4 Termination
-rec succ [^2] This row records the per-step decrease used in the SafeStep proof.
-It does not assert termination of the unrestricted full relation: rec_succ duplicates the step argument $s$ and is the structural source of the full-system termination barrier (cf.
-the conjecture/evidence discussion).
-$\delta$ phase bit (1 $\to$ 0) (n/a) Theorem 4.5 Termination
-eqW refl MPO ($\mu$-first triple) head precedence Theorem 4.6 Termination
-eqW diff MPO ($\mu$-first triple) head precedence Theorem 4.7 Termination
-integrate (delta t) $\mu$ (unique target) (n/a) Theorem 4.8 Termination
-Parenthetical aliases: for substitution convenience simp forms are exposed alongside the main rec lemmas.
+A.5 Auxiliary per-rule orientation (historical)
+Earlier drafts used auxiliary orientation sketches for boundary exploration. They are historical context only and are not the canonical certified path of this paper.
+The canonical per-rule decrease proofs used here are the computable lemmas in Meta/ComputableMeasure.lean (the drop_R_*_c family and measure_decreases_safe_c), with core SafeStep definitions in Meta/SafeStep_Core.lean.
 A.6 Local-join report (Confluence)
@@
 Root eqW a b, a $\neq$ b localJoin_eqW_ne only diff applies
-Root eqW a a, guard ne localJoin_eqW_refl_guard_ne refl blocked; only diff
-Ctx merge a b (n-V, n-E) localJoin_ctx_merge_no_void_neq from root ‘merge’ lemma
+Root eqW a a, guard ne localJoin_eqW_refl_guard_ne both refl and diff blocked (vacuous join)
+Ctx merge a b (n-V, n-E) localJoin_ctx_merge_no_void_neq from root merge lemma
 Ctx eqW a b, a $\neq$ b localJoin_ctx_eqW_ne only diff applies
-Ctx eqW a a, guard localJoin_ctx_eqW_refl_guard_ne refl blocked
-Ctx eqW a a (norm) localJoin_ctx_eqW_refl_if_norm via normalizeSafe
+Ctx eqW a a, guard localJoin_ctx_eqW_refl_guard_ne both refl and diff blocked (vacuous join)
+[…]
 Appendix B Orientation Strategy
 This summarizes the exact orientation strategy realized in the proofs:
-  • Non-duplicating rules use a $\mu$-first lex measure: triples $(\mu,\mathrm{sizeMPO},\delta)$ decrease directly via lemmas.
-  • The rule rec_zero uses a DM multiset drop on $\kappa^{M}$.
-  • The rule rec_succ drops the $\delta$ phase bit (1 $\to$ 0).
-  • Merge-cancel decreases via DM on $\kappa^{M}$ under a base precedence where the redex strictly dominates its pieces.
+  • All guarded rules are oriented in one computable lex order Lex3c over $(\delta\text{-flag},\kappa^{M},\tau)$.
+  • The rule rec_succ is discharged by the outer $\delta\text{-flag}$ drop (1 $\to$ 0).
+  • Multiset-sensitive branches (notably rec_zero) are discharged by DM drops on $\kappa^{M}$, while merge_cancel is discharged by $\tau$.
+  • Tie branches are discharged by strict inequalities in the computable natural rank $\tau$.
+  • Strong normalization follows from well-foundedness of Lex3c via inverse-image and yields wf_SafeStepRev_c.

### v5

@@
 We present KO7, a concrete PRC consisting of seven constructors and eight reduction rules, designed to investigate the foundations of self-verifying computation.
-Standard termination methods fail on this calculus due to a core structural barrier: the duplication of the step argument in the recursor.
-The rec-succ rule redistributes its argument into both the function application and the recursive call ($s\to f(s,\dots)$), an operation that algebraically defeats additive measures ($M_{\text{after}}\geq M_{\text{before}}$ whenever $M_{s}\geq 1$).
+Standard termination methods fail on this calculus due to a core structural barrier: duplication of the step argument in the recursor.
+The rec-succ rule redistributes its step argument into both the function application and the recursive call ($s\to f(s,\dots)$), which blocks strict decrease for additive families (formalized in Section 4).
 Because of this duplication, thirteen distinct internal proof strategies succumb to the calculus’s complexity: additive counters, simple lexicographic measures, polynomial interpretations, single-bit flags, pure ordinal measures, structural depth counters, naive multiset orderings, local helper lemmas, complex hybrid measures, unchecked recursion, pure precedence orders, linear weight functions (KBO-style), and standard tree depth.
@@
 For this fragment, a triple-lexicographic measure (combining phase bits, Dershowitz-Manna multiset ordering, and a computable natural rank) succeeds, supporting a mechanically verified normalizer with proven totality and soundness.
-We state a central conjecture: For Pure Recursive Calculi (PRC), no internally definable method currently known to us proves full-calculus termination.
-Unlike Turing machines (which require an external tape), lambda calculus (which requires variable binding), and recursive functions (which import zero and successor as external primitives), a PRC has none of these external scaffolds.
-On this framing, the termination barrier is positioned at a more primitive layer than those classical settings.
-These findings motivate a form of Pre-Arithmetic Incompleteness.
-Unlike Gödel’s theorems, which require arithmetic sufficient for encoding (Robinson Q), the barrier studied here appears at a more primitive layer: in operator calculi possessing only order and recursion, before arithmetic emerges.
-In this setting, the difficulty of proving full termination from internal methods parallels halting-style limitations at the substrate of operator definitions.
+We state a central conjecture (Section 9): For Pure Recursive Calculi, no internally definable method currently known to us proves full-calculus termination.
+The formally verified impossibility results and the conjecture boundary are detailed in Sections 7 and 9.
 The complete Lean 4 formalization ($\sim$6,000 LOC) is available at https://github.com/MosesRahnama/OperatorKO7.
@@
 There are no binders, types, external axioms, or semantic predicates; the rules are the semantics.
-The primary goals are: (i) a clean proof of strong normalization (SN) that handles duplication; (ii) a certified normalizer that always returns a normal form; (iii) unique normal forms via Newman’s Lemma, with the required local-confluence hypothesis discharged in the artifact; and (iv) a conjecture regarding the unprovability of termination for Pure Recursive Calculi (PRC) using internal methods.
+The primary goals are: (i) a clean proof of strong normalization (SN) that handles duplication; (ii) a certified normalizer that always returns a normal form; (iii) unique normal forms via Newman’s Lemma, with the required local-confluence hypothesis discharged in the artifact; and (iv) a conjecture regarding the limits of internal termination proving for Pure Recursive Calculi (PRC).
 Scope.
@@
 A second conceptual goal is to situate KO7 against results about fixed-target reachability in terminating calculi.
-If an internal provability predicate is defined by “$t$ reduces to $\top$”, then under SN the set decidability of $\{t\mid t\to^{*}\top\}$ is reducible to normalization.
+If an internal provability predicate is defined by “$t$ reduces to a target constant cc”, then under SN the decidability of $\{t\mid t\to^{*}c\}$ is reducible to normalization.
 With confluence, decision reduces to normal-form equality.
 Contributions
-Strong normalization is proven for a guarded fragment of KO7 using a triple-lexicographic measure that combines a phase bit, multiset ordering (Dershowitz-Manna), and a computable natural rank.
-This result supports a certified normalizer with proven totality and soundness.
-Furthermore, a Newman engine for the safe fragment is formalized, and the global local-join hypothesis is fully discharged, yielding unconditional confluence and unique normal forms for the safe fragment.
-Furthermore, we apply established proof theory to classify the boundaries of internal termination proving.
-Drawing on the work of Ferreira and Zantema [8], Buchholz [3], Gallier [9], and Rathjen–Weiermann [12], we observe that classical simplification orders (like LPO and MPO) fall outside our internal-definability boundary for pure calculi.
-These methods achieve termination by introducing the universal Subterm Property ($f(t)>t$) as a precondition, which in proof-theoretic analyses is tied to Kruskal-style well-foundedness arguments and transfinite induction strength (e.g., $\Pi^{1}_{1}$-comprehension).
-Because these orders import external structural axioms and well-ordering principles, they are outside the internal method class used in this paper.
-While not a proof of general non-termination, the work includes formally verified impossibility results for fourteen distinct strategy families under duplication.
-These machine-checked barriers provide evidence of structural limits in internal termination proving for self-referential calculi, directly motivating our central conjecture.
-Highlights (formalization summary)
+The main contributions, formally verified in the artifact, are:
   • SN (SafeStep) via triple lex: Formally proven using a lexicographic measure combining a $\delta$-phase bit, a Dershowitz-Manna multiset rank, and a computable natural rank.
@@
   • Impossibility results: Twelve distinct impossibility barriers (including additive, polynomial, weight-based, and precedence-based measures) are formally machine-checked in Impossibility_Lemmas.lean and Conjecture_Boundary.lean, alongside two meta-theoretical barriers.
+We additionally apply established proof theory to classify the boundaries of internal termination proving.
+Drawing on Ferreira and Zantema [8], Buchholz [3], Gallier [9], and Rathjen–Weiermann [12], we observe that classical simplification orders (like LPO and MPO) fall outside our internal-definability boundary for pure calculi: they achieve termination by importing the universal Subterm Property ($f(t)>t$), which in proof-theoretic analyses is tied to Kruskal-style well-foundedness arguments and transfinite induction strength (e.g., $\Pi^{1}_{1}$-comprehension).
+These machine-checked barriers provide evidence of structural limits in internal termination proving for self-referential calculi, directly motivating our central conjecture.
 2 Background: Term Rewriting, SN, reachability, and Newman
@@
 Termination under duplication typically requires orders beyond plain sizes; a naive size can increase under duplicating rules.
-The cure is a multiset extension of a base order [5] or (recursive) path orders [7].
+In classical rewriting, common responses are a multiset extension of a base order [5] and (recursive) path orders [7]; under our internal-method boundary (Definition 9.1), only the former is admitted after computable projection to numeric payloads, while path orders are treated as external meta-theoretic methods.
 3 The KO7 calculus
@@
 Table 1: KO7’s 8 kernel rules.
-In the full kernel relation, R_eq_refl and R_eq_diff overlap at eqW a a.
-The certified SafeStep fragment adds guards: R_eq_diff requires $a\neq b$, R_merge_void_left/right require $\delta$-flag tie conditions, R_merge_cancel requires both $\delta$-flag and $\kappa^{M}$ guards, R_rec_zero requires a $\delta$-flag guard on $b$, and R_eq_refl requires $\kappa^{M}(a)=0$.
-R_merge_cancel is collapsing.
-(* Note: R_rec_succ duplicates the step argument $s$ (one occurrence on the LHS, two on the RHS).
-This duplication is the termination stressor: the paper proves strong normalization for the guarded SafeStep fragment, not for the unrestricted full relation.)
+In full Step, R_eq_refl and R_eq_diff overlap at eqW a a.
+The SafeStep fragment adds per-rule guards ($\delta$-flag, $\kappa^{M}$, and $a\neq b$ conditions) to resolve this overlap and ensure termination.
+R_rec_succ (*) duplicates the step argument $s$; this duplication is the central termination stressor.
 Rule Head Arity Shape Dup?
@@
 $t::=\mathrm{void}\mid\delta\,t\mid\mathrm{integrate}\,t\mid\mathrm{merge}\,t\,t\mid\mathrm{app}\,t\,t\mid\mathrm{rec}\Delta\,t\,t\,t\mid\mathrm{eqW}\,t\,t.$
-The full kernel relation Step is generated by the 8 rules in Table 1, while the certified artifact is proved for a guarded subrelation SafeStep.
+The full kernel relation Step is generated by the 8 rules in Table 1.
+The certified artifact is proved for a guarded subrelation SafeStep $\subseteq$ Step, which adds per-rule side conditions ($\delta$-flag, $\kappa^{M}$, and disequality guards) that restrict when each rule may fire.
 Tiny example (trace consequences)
@@
   • Strict Syntactic Equality:
-    • If $a$ and $b$ are syntactically identical ($a=b$), then $\mathrm{eqW}\,a\,b\Rightarrow\mathrm{void}$ (provided $\kappa^{M}(a)=0$).
+    • If $a$ and $b$ are syntactically identical ($a=b$), then $\mathrm{eqW}\,a\,b\Rightarrow\mathrm{void}$ (in the safe fragment, provided $\kappa^{M}(a)=0$; unconditional in full Step).
     • If $a\neq b$, then $\mathrm{eqW}\,a\,b\Rightarrow\mathrm{integrate}(\mathrm{merge}\,a\,b)$.
@@
 The guarded relation SafeStep for KO7 is strongly normalizing.
-Genealogy of failures (why DM/MPO)
+Genealogy of failures (why DM and guarded lex components)
 Minimal counterpatterns that motivate the multiset/path components are recorded:
@@
   • Additive bumps on $\kappa$: ties persist on duplicators (Lemma 4.2).
-  • Additive counters $\rho$: by the identity $\rho(\text{after})=\rho(\text{before})-1+\rho(S)$ there is no strict drop when $\rho(S)\geq 1$.
+  • Additive counters $\rho$: by the duplication stress identity above, there is no strict drop when $\rho(S)\geq 1$.
   • Polynomial interpretations that rely on imported arithmetic or hard-coded constants: the duplicating rules force case splits that cannot be oriented without stepping outside the operator language, so these methods fail within the allowed internal toolbox.
@@
 5 A certified normalizer
-A normalization function is defined by well-founded recursion on SafeStepRev, with well-foundedness certified by the computable measure stack $\mu^{3}_{c}$.
-The normalizer operates on the root step relation SafeStep (not the context closure SafeStepCtx); a term is in normal form when no root safe rule applies ($\mathrm{NormalFormSafe}(t)\iff\neg\exists u.\;\mathrm{SafeStep}\;t\;u$).
-Properties proven in the formalization include:
+[…]
 (Totality) $\forall t\,\exists n.~\mathrm{SafeStepStar}(t,n)\land\mathrm{NormalFormSafe}(n)$
@@
 The module also provides corollaries for uniqueness of normal forms and equality of normalizers under star.
-Scope and Guarantees (formalization-accurate)
-The formal results concern a guarded safe subrelation SafeStep for which we prove:
-  • SN (SafeStep). The computable KO7 triple measure $\mu^{3}_{c}$ strictly drops on every SafeStep. This yields a certified normalizer that is total and sound for the safe fragment.
-  • Local-Join Library (SafeStep).
-Local-join lemmas per root shape and context wrappers are provided; the global hypothesis $\mathsf{locAll}$ is discharged in the artifact (locAll_safe), yielding unconditional confluence and unique NFs.
 Critical-pair coverage
@@
   1. 1.
-    Additive Counters: $\kappa(t)+k$ fails for any fixed $k$. The duplication rule $M(\text{after})=M(\text{before})-1+M(s)$ implies no strict drop when $M(s)\geq 1$.
+    Additive Counters: $\kappa(t)+k$ fails for any fixed $k$. The duplicating recursor branch blocks strict additive decrease (Section 4).
   2. 2.
@@
 While full LPO is a classical external method that could theoretically orient such systems, doing so imports this property.
-As argued in Section 8.2, this is a meta-theoretical intervention (via Kruskal-style assumptions) rather than an internally definable proof under Definition 9.1.
+As discussed in Section 9, this is a meta-theoretical intervention (via Kruskal-style assumptions) rather than an internally definable proof under Definition 9.1.
 The code demonstrates pressure toward such an axiom; the text analyzes its scope.
@@
 In a strongly normalizing and confluent calculus, the set $\{t\mid t\Rightarrow^{*}c\}$ for any constant $c$ is decidable via normalization.
-This connects the work to classical results: encoding undecidable properties through reduction to constants violates known theoretical limits.
-This motivates stratified approaches in proof assistants where object-level and meta-level reasoning are kept separate.
-Assumptions (model)
-The model assumes a finite first-order rewrite calculus over finite terms.
-The decision procedure computes a normal form (by SN) and checks whether it is $\top$; with confluence (hence unique normal forms), this reduces to one normalization and an equality test.
-Complexity context
-Small-term reachability in terminating calculi ranges from NP (length-reducing) to NExpTime/N2ExpTime (polynomial interpretations) and PSPACE (KBO), with confluence lowering the length-reducing class to P [1].
-This situates our “normalize and compare” decision procedure for KO7 within the known complexity results.
-Moreover, decidability can be modular for disjoint unions under left-linearity/non-collapsing assumptions [4], while termination alone does not guarantee decidability: even flat non-linear calculi have undecidable reachability/joinability/confluence [10].
+For SafeStep, this is immediate from Cor.
+4.4 and Theorem 6.1: compute $\mathrm{normalizeSafe}(t)$ and compare it with the target constant.
+This section is included only to record that reduction from reachability to normalization is standard in the terminating+confluent setting.
+Complexity bounds for related classes are surveyed in [1, 4, 10].
 9 Conjecture: Full Termination Limits in Pure Recursive Calculi (PRC)
+Sections 47 contain the proved technical core of this paper. The present section states a scoped conjectural boundary motivated by those formal results.
 Sequential Computation
@@
 It applies the “step” term $s$ repeatedly, $n$ times.
-This is functionally identical to the recursor in Gödel’s System T or the step function of primitive recursive arithmetic.
+This is structurally analogous to the recursor in Gödel’s System T, though KO7’s recursor is untyped and unrestricted.
 Both mechanisms allow a calculus to execute a sequence of operations of unbounded length.
@@
 This duplication is the structural barrier.
-As demonstrated in the Impossibility Results (Section 7) and formally witnessed in the artifact by the theorem rec_succ_additive_barrier, this redistribution of $s$ defeats additive termination measures ($M_{\text{after}}\geq M_{\text{before}}$ whenever $M_{s}\geq 1$).
-A calculus that internalizes control flow in this way faces a sharp difficulty when proving full termination with simple internal metrics, because the metric must account for the unbounded potential of the duplicated step function.
+As demonstrated in Section 7 and formally witnessed by rec_succ_additive_barrier, this redistribution defeats strict additive decrease, so simple internal metrics fail to orient the unrestricted branch.
 Escaping the Trap: Meta-Layers vs. Internal Proofs
@@
 In this PRC setting, ordered computation is represented with a recursor.
-The recursor redistributes its step argument across recursive calls.
-This redistribution defeats additive termination measures: $M(\text{after})=M(\text{before})-1+M(s)$, yielding no strict drop when $M(s)\geq 1$.
-Definition 9.1 (Internally definable measure).
-An internally definable termination method for signature $\Sigma$ uses only: DM-multiset lifts of $\mathbb{N}$-valued ranks, and algebraic interpretations definable without importing external axioms.
-No encodings into external arithmetic, no borrowed logic, and no hard-coded constants outside the operator language.
-Importantly, an internally definable method must be derived strictly from the operational semantics (the rewrite rules).
-It cannot inject universal structural axioms from the meta-theory.
-For example, standard simplification orders (LPO/RPO/MPO) are explicitly excluded because they rely on the universal Subterm Property ($f(t)>t$).
-As demonstrated by Ferreira and Zantema [8], this property acts as an imported precondition axiom, artificially establishing a fixed point (an anchor) that the pure calculus does not possess.
-Furthermore, as established in the proof-theoretic literature (e.g., Buchholz [3], Dershowitz & Okada [6]), the well-foundedness of these orders relies on Kruskal’s Tree Theorem, which Gallier [9] notes is a classic independence phenomenon requiring transfinite induction at or above $\varepsilon_{0}$.
-The distinction is: an internally definable method must be self-contained within the operator language; any proof that borrows structural axioms or well-ordering strength from an external theory is performing meta-level reasoning, not internal reasoning.
+The recursor redistributes its step argument across recursive calls, yielding the duplication non-drop mechanism formalized in Section 4.
+Definition 9.1 (Internally Definable Method (design choice; testable)).
+For a finite signature $\Sigma$ and rewrite relation $R$, this paper calls a termination method internally definable iff all of the following checks are met:
+  1. 1.
+    Computable payload extraction: a computable map $\nu:\mathrm{Terms}(\Sigma)\to\mathbb{N}\times(\mathrm{Multiset}\;\mathbb{N}\times\mathbb{N})$ is provided.
+  2. 2.
+    Admitted target order: decrease is measured only in the computable lexicographic order
+[…]
+  3. 3.
+    Per-rule certificate: for every root rule instance $t\to_{R}t^{\prime}$, one proves $\nu(t^{\prime})<_{\mathrm{Lex}}\nu(t)$.
+  4. 4.
+    No imported universal constructor axioms: no global assumptions of the form “for all constructors $f$, f(u)>uf(u)>u” are used unless derivable from $R$ itself.
+This boundary is an explicit methodological design choice for the conjecture scope in this paper. It defines which proof families are counted as “internal” in the audited comparison class.
+[…]
 Definition 9.2 (Pure Recursive Calculus (PRC)).
@@
 Systems of primitive recursive functions fail condition 3 (zero and successor are imported as external primitives, not operator-defined).
-In the sense of this definition, the class targets a more primitive setting than these classical computational models: the termination barrier identified here arises before arithmetic, variable binding, or external memory enter the picture.
+In this scoped definition, the class isolates operator-level recursion without imported arithmetic or external memory.
 9.2 The Subterm Axiom and Artificial Fixed Points
@@
 Thus, in this framework, orienting a PRC with LPO is treated as meta-level reasoning rather than an internally definable termination proof.
-9.3 The Pre-Arithmetic Incompleteness Hierarchy
-The conjecture situates KO7 within a three-layer hierarchy of incompleteness results, presented as requiring progressively more external structure:
-Table 3: Three layers of computational incompleteness.
-PRC (Layer 0) requires only an object-level recursor and no external scaffolding; Turing machines (Layer 1) add an external tape and meta-level iteration; Gödel systems (Layer 2) additionally require arithmetic encoding.
-Layer Calculus class Minimum requirement Barrier
-0 (this work) PRC Order + object-level recursor No known internally definable proof of full termination
-1 (Turing) General computation External tape + meta-level transition Cannot decide halting (undecidable)
-2 (Gödel) Arithmetic calculus Robinson Q (++ and $\times$) Cannot prove own consistency internally
-Why Layer 0 Is Placed Below Layer 1
-A classical Turing machine performs sequential computation via a meta-level transition function $\delta:Q\times\Gamma\to Q\times\Gamma\times\{L,R\}$.
-The iteration (“run for $k$ steps”) is defined outside the machine as a sequence of configuration yields.
-The tape is external memory; there is no operator within the term language that internalizes the step.
-In a PRC, by contrast, the recursor $\mathrm{rec}\Delta$ is an object-level constructor: “do the next step” is a rewrite rule applied from inside the term, not a meta-operation on a configuration.
-This is why a Turing machine fails condition 2 of Definition 9.2 and falls outside the PRC class.
-Why Layer 0 Is Placed Below Layer 2
-Gödel’s incompleteness theorems require arithmetic sufficient for encoding, at minimum Robinson arithmetic $Q$, with addition, multiplication, zero, and successor as imported primitives.
-Presburger arithmetic (addition only) and Skolem arithmetic (multiplication only) are both decidable and complete; they escape Gödel because they lack the encoding power of full $Q$.
-A PRC requires none of this: no imported arithmetic, no Gödel numbering, no self-referential encoding.
-The termination barrier at Layer 0 arises from structural duplication in the recursor (a constructive phenomenon) rather than from a diagonal self-reference argument.
-Arithmetic itself is a higher-layer phenomenon: it emerges when the PRC operators are later interpreted (e.g., $\delta$ as successor, $\mathrm{rec}$ as primitive recursion).
-The Layer 0 barrier exists before that interpretation is imposed.
-This explains why single-level Gödel encodings cannot coexist with globally terminating proof search; stratification is required.
+9.3 Layered perspective (discussion only)
+For context, we include a layered comparison with Turing-style and arithmetic incompleteness frameworks in Appendix A.2.
+This comparison is intentionally conjectural discussion, not a theorem-level reduction or embedding result.
 Conjecture 9.3 (Scoped Full-Termination Conjecture for Pure Recursive Calculi).
@@
   • Additive and polynomial methods are formally excluded in the audited family.
-The redistribution identity $M(\text{after})=M(\text{before})-1+M(s)$ yields no strict drop when $M(s)\geq 1$; polynomial interpretations without imported arithmetic degrade to additive measures and fail by the same identity (Lemma 4.2).
+By the duplication non-drop mechanism from Section 4, additive counters do not strictly decrease on unrestricted duplicating recursors; polynomial interpretations without imported arithmetic collapse to the same obstruction.
 Machine-checked barriers in Meta/Conjecture_Boundary.lean formally witness failure of additive counters, simple lexicographic pairs, flag-only methods, polynomial multiplicative ties, structural depth, and constellation-style hybrids (Section 7).
@@
 Under Definition 9.1, importing this axiom violates the “no imported axioms” boundary for PRCs; therefore, such arguments are treated as meta-theoretical rather than internal proofs.
-    Note on orientation layers: The canonical SafeStep termination proof in this paper is the computable stack $(\delta,\kappa^{M},\tau)$ from Meta/ComputableMeasure.lean: rec_zero is discharged by DM on $\kappa^{M}$, rec_succ by the $\delta$ drop, and merge_cancel/eqW tie branches by $\tau$.
-MPO/LPO appear only as external meta-theoretic comparators in the boundary discussion, not as the internal certified certificate.
+    Note: The canonical SafeStep certificate is the computable stack $(\delta,\kappa^{M},\tau)$ from Meta/ComputableMeasure.lean (Section 4); MPO/LPO appear only as external comparators in the boundary discussion.
 Any proof that relies on external axioms, imported arithmetic, or meta-level encodings is outside this scope.
@@
 Evidence
-In KO7, the rec-succ rule has the critical shape: the step argument $s$ appears twice on the RHS, and when $s$ contains nested redexes, additive measures fail to decrease strictly.
-The redistribution identity $M(\text{after})=M(\text{before})-1+M(s)$ yields no strict drop when $M(s)\geq 1$.
-The guarded SafeStep fragment terminates because the $\delta\text{-flag}$ restricts allowable step arguments; without such restriction, the full calculus admits unbounded recursive depth.
-At present, no internally definable ranking-function closure is known in this setting without external methods (axioms, encodings, or borrowed arithmetic).
+In KO7, the rec-succ rule has the critical shape: the step argument $s$ appears twice on the RHS, and when $s$ contains nested redexes, additive measures fail to decrease strictly (Section 4).
+The guarded SafeStep fragment terminates because the $\delta$-phase and $\kappa^{M}$ guards restrict this branch; without such restriction, no internally definable ranking-function closure is currently known to us without external methods (axioms, encodings, or borrowed arithmetic).
 10 Formalization structure
@@
   • Impossibility results: Contains the verified counterexamples for additive measures and the proofs that simple measures fail under duplication.
-All theorem statements currently asserted in the build are formally proven without unproven postulates.
+All theorem statements currently asserted in the build are formally proven without sorry, admit, or unsafe declarations.
 Confluence theorems in Meta/Newman_Safe.lean are stated with a global local-join parameter ($\mathsf{locAll}$).
-The artifact discharges this parameter via locAll_safe from Meta/Confluence_Safe.lean to obtain unconditional confluentSafe and unique normal forms for SafeStep.
+The artifact discharges this parameter via localJoin_all_safe from Meta/Confluence_Safe.lean, yielding locAll_safe in Meta/Newman_Safe.lean, to obtain unconditional confluentSafe and unique normal forms for SafeStep.
 The formal development (Lean 4, $\sim$6,000 LOC; dual-licensed for academic and commercial use) is available at https://github.com/MosesRahnama/OperatorKO7 and includes build instructions and the toolchain version.
@@
 An explicit non-local-join witness at $\mathrm{eqW}\;\mathrm{void}\;\mathrm{void}$ is established for the full relation, motivating the SafeStep fragment.
-The decidability result explains the clash between single-level Gödel encodings and global termination.
-The impossibility of additive bumps under duplication is formally verified.
+Impossibility results for twelve distinct strategy families (including additive, polynomial, weight-based, and precedence-based measures) are formally verified under duplication, and this evidence motivates a scoped conjectural boundary for unrestricted PRC termination under Definition 9.1.
 Appendix A Addendum
@@
 The authoritative mechanized kernel is the Lean file OperatorKO7/Kernel.lean in the repository https://github.com/MosesRahnama/OperatorKO7. The safe fragment and proofs are under OperatorKO7/Meta/.
-A.2 Auxiliary orientation sketch (DM/MPO)
+A.2 Layered Perspective (Conjectural Discussion)
+This subsection records a discussion framework used to contextualize the PRC conjecture. It is not a proved reduction between formalisms.
+Table 3: Conjectural layered comparison (discussion only).
+Layer Calculus class Minimum requirement Classical barrier
+0 (scoped conjecture) PRC Order + object-level recursor No known internally definable proof of unrestricted full termination
+1 (Turing) General computation External tape + meta-level transition Halting undecidability
+2 (Gödel) Arithmetic calculus Robinson Q (++ and $\times$) Incompleteness/consistency limits
+Interpretation note
+The table is a framing device: it compares the amount of external structure assumed in each setting.
+It does not claim a theorem stating that Layer 0 is formally reducible to or strictly below Layers 1 or 2.
+A.3 Auxiliary orientation sketch (historical; external MPO comparator)
 This subsection records historical orientation notes used for boundary analysis.
 It is not the canonical SafeStep termination certificate in this paper, which is the computable stack in Meta/ComputableMeasure.lean.
+Any MPO reference below is for external comparison only and is excluded from internally definable methods by Definition 9.1.
 An auxiliary triple $\widetilde{\mu}^{3}(t)=(\delta\text{-flag}(t),\,\kappa^{M}(t),\,r_{\mathrm{aux}}(t))$ is ordered lexicographically:
   • $\delta$-flag: phase bit dropping on the successor recursion branch.
-  • Multiset $\kappa^{M}$: a Dershowitz-Manna multiset extension over a base precedence/status (or MPO head precedence) orienting redexes strictly above pieces; covers duplicators.
+  • Multiset $\kappa^{M}$: a Dershowitz-Manna multiset extension over an internal base numeric order on interpreted payloads, orienting redexes strictly above pieces; covers duplicators.
   • Tie-break rank $r_{\mathrm{aux}}$: resolves non-duplicating ties when the first two coordinates are equal.
@@
 It is emphasized that $\kappa^{M}$ aggregates via multiset operations (union $\cup$ and addition ++), not scalar numeric addition.
-For duplicating rules, the multiset of piece-weights on the RHS is DM-smaller than the singleton multiset containing the LHS redex weight, yielding a strict drop in the $\kappa^{M}$ component.
-For non-duplicating rules, $\kappa^{M}$ ties by definitional equality and $r_{\mathrm{aux}}$ resolves the branch.
-In particular, for unguarded instances of eqW-refl and merge-cancel a $\kappa$-branch strategy is employed: when $\kappa^{M}(a)\neq 0$ a left-lex drop is obtained via DM; when $\kappa^{M}(a)=0$ the $\kappa^{M}$ component ties by rfl and the strict decrease is witnessed in the $r_{\mathrm{aux}}$ coordinate (right branch of Prod.Lex).
+For duplicating rules, the multiset of piece-weights on the RHS is DM-smaller than the singleton multiset containing the LHS redex weight, yielding a strict drop in the $\kappa^{M}$ component where applicable.
+For non-duplicating rules, $\kappa^{M}$ may tie by definitional equality and $r_{\mathrm{aux}}$ resolves the branch.
+In the canonical SafeStep certificate, merge_cancel is guarded by $\kappa^{M}(t)=0$, so strict decrease is discharged in the $r_{\mathrm{aux}}$$\tau$ coordinate.
+For unguarded eqW-refl analysis, a $\kappa$-branch split is used: when $\kappa^{M}(a)\neq 0$, a left-lex DM drop applies; when $\kappa^{M}(a)=0$, $\kappa^{M}$ ties by rfl and strict decrease is witnessed in the $r_{\mathrm{aux}}$ coordinate (right branch of Prod.Lex).
 Short witness snippets (toy duplication)
@@
 Use $X=\varnothing$, $Y=\{\mathrm{size}(x),\mathrm{size}(y)\}$, $Z=\{S(x,y)\}$ to conclude $Y<_{\mathrm{DM}}Z$.
-MPO triple weight.
+Historical MPO comparator (external).
 $\mathrm{weight}(\mathrm{pair}\,a\,b)=(\mathrm{headRank}(a),\,\mathrm{size}(a),\,\mathrm{size}(b))$ with $\mathrm{headRank}(\mathrm{s}\,\_)=2$, else $1$.
@@
 Figure 1: Auxiliary triple-lex components. Duplicators decrease via $\kappa^{M}$; non-duplicating ties via $r_{\mathrm{aux}}$.
-A.3 Newman scope (guarded safe relation)
-SN + local confluence implies confluence; in the artifact this is instantiated for the safe relation via an Acc-based star-star join.
-Scope note: all confluence statements and their Newman instantiations are for SafeStep only.
-Full Step has an explicit non-local-join witness at root, eqW void void; accordingly full-step confluence is not claimed.
 A.4 Module map
-The verification logic is distributed across key modules for SN, normalizer, and confluence engine.
-Local confluence lemmas are proved separately.
-Note: context closure and local-join lemmas are explicitly instantiated for the SafeStep fragment (SafeStepCtx) rather than the full relation.
+The verification logic is distributed across the following modules:
+  • Kernel.lean — 7-constructor Trace inductive, 8-rule Step relation, reflexive-transitive closure Star.
+  • Meta/SafeStep_Core.leanweight, kappaM, deltaFlag, guarded SafeStep relation.
+  • Meta/ComputableMeasure.leantau, Lex3c, per-rule decrease lemmas (drop_R_*_c), wf_SafeStepRev_c.
+  • Meta/Normalize_Safe.lean — Certified normalizer via WellFounded.fix; totality and soundness proofs.
+  • Meta/SafeStep_Ctx.lean — Context closure SafeStepCtx for the safe fragment.
+  • Meta/Confluence_Safe.lean — Local-join lemmas per root/context shape; localJoin_all_safe.
+  • Meta/Newman_Safe.lean — Newman’s lemma instantiation; confluentSafe, unique normal forms.
+  • Meta/Impossibility_Lemmas.lean — Machine-checked failed-measure catalog.
+  • Meta/Conjecture_Boundary.lean — No-go theorems, GlobalOrients framework.
+  • Meta/Operational_Incompleteness.lean
+    Probe predicates (P1–P3), InternallyDefinableMeasure.
+Context closure and local-join lemmas are explicitly instantiated for the SafeStep fragment rather than the full relation.
 A.5 Auxiliary per-rule orientation (historical)
@@
 […]
-Appendix B Orientation Strategy
-This summarizes the exact orientation strategy realized in the proofs:
-  • All guarded rules are oriented in one computable lex order Lex3c over $(\delta\text{-flag},\kappa^{M},\tau)$.
-  • The rule rec_succ is discharged by the outer $\delta\text{-flag}$ drop (1 $\to$ 0).
-  • Multiset-sensitive branches (notably rec_zero) are discharged by DM drops on $\kappa^{M}$, while merge_cancel is discharged by $\tau$.
-  • Tie branches are discharged by strict inequalities in the computable natural rank $\tau$.
-  • Strong normalization follows from well-foundedness of Lex3c via inverse-image and yields wf_SafeStepRev_c.